### v1

# Scale-Dependent Multifractality in Bitcoin Realised Volatility: Implications for Rough Volatility Modelling

Scale-Dependent Multifractality in Bitcoin Realised Volatility: Implications for Rough Volatility Modelling
Milan Pontiggia
Master’s Student, MAGEFI – Magistère in International Economics and Finance University of Bordeaux, France
July 1, 2025
We assess the applicability of rough volatility models to Bitcoin realised volatility using the normalised $p$-variation framework of Cont and Das, (2024). Applying this model-free estimator to high-frequency Bitcoin data from 2017 to 2024 across multiple sampling resolutions, we find that the normalised statistic remains strictly negative throughout, precluding the estimation of a valid roughness index. Stationarity tests and robustness checks reveal no significant evidence of non-stationarity or structural breaks as explanatory factors. Instead, convergent evidence from three complementary diagnostics, namely Multifractal Detrended Fluctuation Analysis, log-log moment scaling, and wavelet leaders, reveals a multifractal structure in Bitcoin volatility. This scale-dependent behaviour violates the homogeneity assumptions underlying rough volatility estimation and accounts for the estimator’s systematic failure. These findings suggest that while rough volatility models perform well in traditional markets, they are structurally misaligned with the empirical features of Bitcoin volatility.
1 Introduction
1.1 Background and Motivation
The rough volatility framework has emerged as a prominent approach for modelling the fine-scale dynamics of financial market volatility. This development is grounded in empirical findings that indicate that realised volatility paths possess low Hölder regularity, which implies irregularity in their sample paths at high frequencies. Rough stochastic volatility models have been shown to replicate empirical features commonly observed in financial markets, including volatility clustering, short-term persistence, and the skew and term structure of implied volatility surfaces in equity and index markets Gatheral et al., (2018); Fukasawa et al., (2022). These observations have motivated attempts to extend the applicability of rough volatility models to nontraditional asset classes.
However, the extent to which this framework generalises to structurally distinct markets remains an open empirical question. Bitcoin, the most liquid and actively traded cryptocurrency, offers a particularly informative out-of-sample test. Unlike traditional equity or FX markets, Bitcoin exhibits fragmented liquidity across exchanges Makarov and Schoar, (2020), frequent discontinuities due to jumps Lahmiri and Bekiros, (2020), regime-switching volatility Caporale and Zekokh, (2021), and an evolving market microstructure shaped by protocol changes and shifting incentives Easley et al., (2019). This motivates a closer empirical examination of whether rough volatility assumptions hold in such a market.
In this context, the term roughness refers to the local irregularity of the volatility path, quantified in terms of its Hölder regularity. A process is said to be rough if its sample paths exhibit lower regularity than the sample paths of Brownian motion. This notion of roughness is mathematically formalised via the concept of a roughness index, the reciprocal of the smallest $p\geq 1$ such that the process admits finite $p$-variation along a refining partition sequence. In rough volatility models, $H$ captures the scaling properties of volatility at infinitesimal time scales, and accurately estimating it is essential for determining whether volatility exhibits statistically significant roughness.
Prior applications of rough volatility methodologies to Bitcoin have predominantly employed parametric inference techniques or heuristic scaling diagnostics, often without systematically assessing assumptions of homogeneous Hölder continuity Bianchi et al., (2022). Empirical studies on cryptocurrencies more broadly, including Liu and Tsyvinski, (2021), have concentrated on asset pricing and factor models, without directly addressing the scaling properties of volatility. The normalised $p$-variation estimator recently introduced by Cont and Das, (2024) provides a model-free approach to estimating the pathwise roughness of volatility signals. This estimator offers two notable advantages over traditional methods based on log-volatility increments or variogram regressions: (i) it exhibits improved asymptotic properties under general conditions, and (ii) it is less sensitive to microstructure noise and finite-sample distortions. These properties make it especially suitable for high-frequency volatility data in contexts like cryptocurrency markets, where noise and irregular sampling are common.
This study builds on the normalised $p$-variation estimator introduced by Cont and Das, (2024) to perform a systematic evaluation of the roughness of Bitcoin realised volatility across a range of sampling frequencies and time horizons. This analysis examines the applicability of rough volatility models to Bitcoin and contributes to a deeper understanding of the structural features that differentiate cryptocurrency volatility dynamics.
1.2 Research Gap
While the rough volatility literature has made significant progress in recent years, its empirical evidence is drawn primarily from equity and index markets. In these settings, various empirical methodologies, including parametric estimation and $p$-variation diagnostics, have provided evidence for the existence of a global scaling regime consistent with rough stochastic volatility models. However, the applicability of such models to structurally distinct markets such as cryptocurrencies remains largely unexplored.
Bitcoin’s volatility displays structural features such as multiscale clustering, episodic jumps, and irregular liquidity that differ markedly from the conditions under which rough volatility models have been validated.For instance, Bianchi et al. (2022) report low Hurst exponents for Bitcoin by fitting rough stochastic volatility models and performing scaling regressions on realised volatility. However, their approach implicitly assumes a global scaling regime and stationarity, without subjecting these conditions to empirical verification. These assumptions are critical, and their violation under multifractal dynamics or structural irregularities undermines the reliability of such roughness estimates. Crucially, no prior work has systematically examined whether Bitcoin realised volatility meets the homogeneity and stationarity conditions required for consistent estimation via the normalised $p$-variation method.
Moreover, recent research indicates that the observed roughness of realised volatility may arise from estimation error rather than genuine features of the spot volatility process Cont and Das, (2024); Rogers, (2019). In the case of Bitcoin, where multifractality and non-stationarity are theoretically plausible features, this raises serious doubts about the robustness and interpretive value of roughness estimates.
To our knowledge, no prior empirical study has systematically applied the $p$-variation diagnostic to Bitcoin realised volatility, nor has tested its structural compatibility with the assumptions underlying rough volatility models. It remains empirically unestablished whether Bitcoin volatility admits a well-defined roughness exponent. This study addresses this gap by applying the model-free normalised $p$-variation estimator alongside complementary diagnostic methods.
1.3 Research Question and Empirical Puzzle
This study tests whether Bitcoin realised volatility satisfies the homogeneous scaling assumptions necessary for consistent roughness estimation under the rough volatility paradigm.
We investigate this question by applying the normalised $p$-variation methodology of Cont and Das, (2024) to Bitcoin realised volatility across multiple years and sampling frequencies. Preliminary results indicate that, in contrast to equity and foreign exchange markets, Bitcoin volatility does not appear to admit a well-defined roughness index. This stands in contrast to prior findings for equity and foreign exchange markets, where roughness estimates have exhibited consistent behaviour and robust interpretability.
This empirical anomaly raises a fundamental question: is the failure of roughness estimation for Bitcoin a finite-sample artefact, a manifestation of market-specific features such as multifractality, or evidence that rough volatility models are structurally incompatible with this asset class? Resolving this question is central to understanding the volatility dynamics of cryptocurrency markets and to evaluating the broader applicability of rough volatility models beyond traditional financial assets.
1.4 Hypotheses
We formulate the following working hypotheses to guide our empirical investigation:
H1. Bitcoin volatility does not admit a global scaling law compatible with the normalised $p$-variation framework:
Bitcoin volatility may not admit a well-defined global scaling law consistent with the assumptions of the normalised $p$-variation framework. Structural features such as microstructure noise, fragmented liquidity, and large volatility jumps can disrupt scaling regularity, making the estimation of a roughness index theoretically ill-posed. Empirically, this would manifest as a $\log W(L,K,p)$ curve that fails to cross zero for any $p$.
H2. Estimation Artefacts and Finite-Sample Effects:
Even if rough scaling holds in principle, practical estimation may fail due to finite-sample distortions, data irregularities, or extreme returns. In this scenario, the systematic negativity of $\log W(L,K,p)$ would indicate a breakdown of the estimator under limited or noisy data, rather than structural violations of scaling. This hypothesis attributes the failure to the method’s operational limitations, not to the underlying process.
H3. Non-Stationarity or Regime Shifts:
Bitcoin volatility may exhibit non-stationary or structural breaks, violating key assumptions of the $p$-variation estimator. Regime changes, persistent trends, or time-varying scaling could compromise the interpretability of the $W(L,K,p)$ statistic. We assess this hypothesis by applying the Augmented Dickey-Fuller (ADF) test for unit roots, computing rolling mean and variance statistics to evaluate local stability, and detecting structural breaks using binary segmentation with an $L_{2}$-based cost function, as detailed in Section 3.6.1.
H4. Multifractality of Bitcoin Volatility:
Bitcoin volatility may exhibit an inherently multifractal temporal structure, where different segments exhibit varying local scaling exponents. This heterogeneity precludes the existence of a single global roughness index and is incompatible with the normalised $p$-variation framework. We investigate this hypothesis using complementary diagnostics: Multifractal Detrended Fluctuation Analysis (MF-DFA), log-log scaling of moments, and Wavelet Leaders Multifractal Analysis.
1.5 Contributions
This study makes three contributions to the literature on volatility modelling and the structural properties of cryptocurrency markets.
First, it offers an extensive application of the normalised $p$-variation framework to Bitcoin realised volatility across multiple years and sampling frequencies, making the analysis more comprehensive than any prior roughness estimation in cryptocurrency markets. The analysis reveals a systematic and robust failure of roughness estimation across all empirical configurations.
Second, we demonstrate that the failure of roughness estimation cannot be attributed to finite-sample effects or non-stationarity. Extensive stationarity tests and window stability analyses exclude estimation artefacts and regime changes. Multifractal diagnostics, namely Multifractal Detrended Fluctuation Analysis, log-log moment scaling, and wavelet leaders analysis, based on established fractal analysis methods Kantelhardt et al., (2002); Wendt et al., (2007), confirm that Bitcoin volatility exhibits a multifractal temporal structure. This structure is incompatible with the homogeneous scaling assumptions required by rough volatility models.
Third, the results contribute to the evaluation of the rough volatility paradigm by showing that, while effective in equity and foreign exchange markets, rough stochastic volatility models face structural challenges when applied to Bitcoin realised volatility. This finding underscores the need for models that explicitly incorporate multifractal and scale-dependent dynamics to capture the unique characteristics of cryptocurrency markets. Examples of such alternatives include the Multifractal Random Walk Bacry et al., (2001) and Multifractal Stochastic Volatility Calvet and Fisher, (2002) models.
2 Literature Review
2.1 Rough Volatility and the Hurst Paradigm
The rough volatility paradigm models financial market volatility as a stochastic process whose sample paths are rougher than those of Brownian motion. Motivated by empirical evidence that volatility exhibits low Hölder regularity at high frequencies, this approach posits that volatility dynamics are well described by fractional processes with Hurst exponents $H<0.5$ Gatheral et al., (2018). In such models, volatility paths display pronounced fine-scale irregularity (indicative of low short-term regularity) but not necessarily long-range dependence. This roughness aligns with the volatility clustering and persistent fluctuations observed in high-frequency equity market data.
Empirical support for the rough volatility hypothesis has emerged from multiple methodological approaches. These include the log-regression of realised volatility increments Gatheral et al., (2018), parametric estimation of fractional stochastic volatility models Fukasawa et al., (2022), and non-parametric diagnostics based on $p$-variation Cont and Das, (2024). Across these methods, estimated Hurst exponents in equity and index markets typically fall in the range 0.1 to 0.3. These findings have motivated the adoption of rough stochastic volatility models, originally proposed by Gatheral and Rosenbaum and empirically validated by Bennedsen, Lunde, and Pakkanen Bennedsen et al., (2016), in both academic research and quantitative finance practice. Rough models have demonstrated empirical success in capturing the fine-scale behaviour of implied volatility surfaces and short-term volatility dynamics.
The applicability of the rough volatility paradigm across asset classes remains contested. Rogers, (2019) and Cont and Das, (2024) have emphasised that low roughness estimates derived from realised volatility may not reflect intrinsic properties of the spot volatility process, but instead result from structural features of the data or estimation artefacts. For example, Rogers, (2019) showed that microstructure noise or irregular sampling can generate artificially low Hurst exponents, cautioning against interpreting realised path roughness as direct evidence of true volatility roughness. These confounding effects, including multifractal scaling, may induce apparent roughness even when the underlying volatility process is not genuinely rough. This underscores the need for model-free diagnostics when extending rough volatility models beyond traditional markets.
2.2 The Cont–Das Normalised $p$-Variation Estimator
To address the limitations of traditional approaches, we adopt the framework introduced by Cont and Das, (2024). This framework provides a model-free, pathwise-consistent estimator of roughness based on the scaling behaviour of the normalised $p$-variation statistic $W(L,K,p)$, computed over a sequence of block partitions of the observed time series. Intuitively, $W(L,K,p)$ measures how the $p$-th variation of the process accumulates across blocks, normalised in such a way that for a self-similar process with a global scaling law, the statistic equals one exactly at the correct scaling exponent $p^{*}$. Thus, plotting $\log W(L,K,p)$ as a function of $1/p$, one expects a zero-crossing at a unique value $p=p^{*}$, which defines the variation index and yields the corresponding roughness index $H=1/p^{*}$. The existence of this zero-crossing serves as a rigorous criterion for determining whether the process admits a global monofractal scaling regime consistent with Hölder-type roughness. For reference, the formal definition of $W(L,K,p)$ is provided in Section 3.1.
The normalised $p$-variation estimator is non-parametric, robust to non-Gaussian features, and directly applicable to realised volatility series. It serves not only as a roughness estimator but also as a structural diagnostic of the underlying process. Specifically, the presence or absence of a zero-crossing in $\log W(L,K,p)$ determines whether the data admit a well-defined global scaling law consistent with the rough volatility paradigm. If no zero-crossing is observed, it implies that no single roughness exponent can characterise the process, indicating a fundamental incompatibility with monofractal scaling assumptions.
Recent applications of the method to equity markets have validated its efficacy Cont and Das, (2024), producing roughness estimates consistent with the existing literature. In particular, Cont and Das apply the normalised $p$-variation estimator to 5-minute realised volatility of the S&P 500 using the Oxford-Man Institute dataset. They observe a well-defined zero-crossing in $\log W(L,K,p)$, with the resulting roughness index estimated in the range $H\in[0.05,0.25]$, consistent with prior findings from log-regression methods.
2.3 Multifractality in Financial Time Series
Multifractality characterises time series in which different segments exhibit heterogeneous local scaling behaviour, thereby precluding the existence of a single global scaling exponent. In such processes, the scaling exponents vary continuously with the moment order, leading to non-linear moment scaling functions or curvature in the wavelet spectrum Bacry et al., (2001); Calvet and Fisher, (2002). This stands in contrast to monofractal processes, which are governed by a single exponent that applies uniformly across all time scales and moments.
Empirical research has established that financial time series often exhibit multifractal properties. Foundational models such as the Multifractal Random Walk Bacry et al., (2001) and the Multifractal Stochastic Volatility model Calvet and Fisher, (2002) account for stylised features such as volatility clustering, heavy tails, and long memory through underlying multifractal dynamics. Subsequent studies have confirmed these properties across equities, foreign exchange, and commodities. More recent work has extended this evidence to cryptocurrencies. High-frequency Bitcoin returns and trading volume exhibit multifractal scaling when analysed using Multifractal Detrended Fluctuation Analysis Mensi et al., (2019); Matteo, (2007), and complex multiscale dependence has been observed in cross-correlation structures Watorek et al., (2022). Additional research shows that Bitcoin realised volatility also displays a multifractal structure, though weaker than that of returns Zhou et al., (2022). These findings support the view that multifractality is a robust and persistent feature of financial markets and offer a plausible explanation for the structural deviations from monofractal scaling observed in Bitcoin volatility.
Multifractality is fundamentally incompatible with the rough volatility paradigm, which assumes homogeneous Hölder regularity across scales. A multifractal process does not admit a single Hurst exponent, its apparent roughness varies with the moment order or temporal scale considered. As a result, roughness estimators based on global scaling laws, such as the normalised $p$-variation estimator, are theoretically ill-posed in this context and may yield misleading conclusions. In such settings, the absence of a well-defined roughness index is not a limitation of the estimator but reflects a structural property of the underlying process Cont and Das, (2024).
3 Data and Methodology
3.1 Mathematical Framework
$p$-Variation and Hölder Regularity:
Let $X=\{X_{t}\}_{t=1}^{L}$ be a real-valued function or time series. Its $p$-variation over the discrete index set $\{1,\dots,L\}$ is defined as
$V_{p}(X)=\sup_{\mathcal{P}}\sum_{i=1}^{n}|X_{t_{i}}-X_{t_{i-1}}|^{p},$ (1)
where the supremum is taken over all finite partitions $\mathcal{P}=\{0=t_{0}<t_{1}<\dots<t_{n}=L\}$. If $V_{p}(X)<\infty$, we say that $X$ has finite $p$-variation.
A related concept is the Hölder exponent $H\in(0,1)$, which characterises the local regularity of $X$ via the condition
$|X_{t}-X_{s}|\leq C|t-s|^{H}\quad\text{for all }s,t\text{ in a neighborhood}.$
Smaller values of $H$ correspond to rougher paths. Brownian motion has $H=0.5$, while rough volatility models typically assume $H\in(0.1,0.4)$. The roughness index we estimate serves as a nonparametric proxy for local Hölder regularity.
Normalised $p$-Variation and Roughness Index:
Let $X=\{X_{t}\}_{t=1}^{L}$ denote a discrete time series with $L$ observations. Given a partition of $X$ into $K$ consecutive, non-overlapping blocks of size $n=L/K$, define the block sums $S_{j}=\sum_{i=(j-1)n+1}^{jn}X_{i}$ for $j=1,\dots,K$.
The normalised $p$-variation statistic is then defined as
$W(L,K,p)=\frac{1}{K}\sum_{j=1}^{K}|S_{j}|^{p}.$ (2)
For a process exhibiting homogeneous scaling, $\log W(L,K,p)$ is expected to be strictly decreasing in $1/p$ and to cross zero at a unique value $p=p^{*}$ such that $W(L,K,p^{*})=1$. The roughness index is then defined as
$\hat{H}=\frac{1}{p^{*}}.$ (3)
This index $\hat{H}$ captures the fine-scale regularity of the path $X$. A value $\hat{H}<0.5$ indicates rougher-than-Brownian behaviour. This construction forms the basis of the normalised $p$-variation estimator of Cont and Das, (2024), which detects roughness by identifying the moment $p^{*}$ at which the aggregated variation normalises. The estimate $\hat{H}=1/p^{*}$ serves as a nonparametric, model-free proxy for the path’s Hölder-type regularity.
Multifractality and Scaling Exponents:
A stochastic process $X=\{X_{t}\}_{t=1}^{L}$ is said to exhibit multifractality if its scaling behaviour varies with the moment order $q$. This is characterised by a non-linear dependence of structure function exponents on $q$.
Log-moment scaling diagnostics evaluate the empirical relation
$M_{q}(\tau)=\mathbb{E}[|X_{t+\tau}-X_{t}|^{q}]\sim\tau^{\zeta_{q}},$ (4)
where $\zeta_{q}$ is the $q$-th order scaling exponent and $\tau$ denotes the time scale. In practice, $\zeta_{q}$ is estimated by regressing $\log M_{q}(\tau)$ on $\log\tau$ over a range of scales. Multifractality is indicated by the non-linearity of the $\zeta_{q}$ curve in $q$, monofractal processes correspond to linear $\zeta_{q}$.
Multifractal detrended fluctuation analysis (MF-DFA) estimates the generalised Hurst exponent $H(q)$ via the relation
$F_{q}(s)\sim s^{H(q)},$ (5)
where $F_{q}(s)$ is the $q$-th order fluctuation function over window size $s$. As with $\zeta_{q}$, curvature in $H(q)$ implies multifractality, while constancy indicates monofractal behaviour.
Wavelet leaders multifractal analysis estimates scaling via the moments of wavelet leader coefficients $d_{j,k}$ at scale $2^{j}$:
$\mathbb{E}[|d_{j,k}|^{q}]\sim 2^{j\zeta_{q}}.$ (6)
Linear regression of $\log_{2}\mathbb{E}[|d_{j,k}|^{q}]$ on $j$ yields $\zeta_{q}$, and the width $\zeta_{\max}-\zeta_{\min}$ serves as a summary statistic for the strength of multifractality. As a scalar diagnostic, we compute the spectrum width defined as $\zeta_{q_{\max}}-\zeta_{q_{\min}}$, typically over $q\in[-4,+4]$. Larger widths indicate stronger multifractal heterogeneity.
Across all methods, multifractality is diagnosed through the non-linearity and spectrum width of $q\mapsto H(q)$ or $q\mapsto\zeta_{q}$, reflecting the heterogeneity of local scaling exponents. These diagnostics are applied over appropriate intermediate scales by selecting intermediate time scales that avoid both microstructure noise (short scales) and non-stationary drift (long scales) (see Kantelhardt et al., (2002) for MF-DFA and Wendt et al., (2007) for wavelet leader analysis). The definitions presented in this section underpin the empirical diagnostics and estimation procedures reported in Sections 3.6.2 and 4.3.
3.2 Data Description and Preprocessing
This study employs a high-frequency dataset of BTC/USD prices recorded at one-minute intervals, derived from the publicly available “Bitcoin Historical Data” dataset on Kaggle. The dataset aggregates minute-level transaction data from the Bitstamp exchange via its public API. The full dataset spans from January 2012 to May 2025 and includes open, high, low, and close prices, along with traded volume in BTC, timestamped in UTC. It is updated daily through an automated pipeline hosted on GitHub Actions, which retrieves and integrates historical data directly from Bitstamp.
Although the dataset includes automated deduplication and merge routines, its documentation acknowledges potential limitations in data completeness and consistency, including gaps due to exchange outages or API availability. No formal data validation is provided. To mitigate these issues, we restrict our analysis to the 2017-2024 period, during which data quality is sufficiently stable for high-frequency volatility estimation.
To ensure consistency and minimise missing data effects, for each calendar year $Y$, we systematically selected the most complete 90-day window of consecutive minute-level BTC/USD observations. We opted to analyse one representative 90-day high-quality segment per year to control for seasonality and to ensure stationarity over each segment. Using longer or multiple windows risked including structural changes or requiring imputation for extensive gaps. Let $T=129{,}600$ minutes denote the target window length. The search space was defined as the set of all candidate windows of length $T$, starting at time $t_{s}$ and ending at $t_{e}=t_{s}+T-1$, with $t_{s}$ incremented in daily steps (1,440 minutes). For each candidate window, we constructed the corresponding sequence of timestamps $\{t_{i}\}_{i=1}^{T}$ and computed the number of missing Close price observations $m(t_{s})$. The optimal window $[t_{s}^{*},t_{e}^{*}]$ was selected as:
$(t_{s}^{*},t_{e}^{*})=\arg\min_{t_{s}}m(t_{s})$
In the event of ties, where multiple windows contained the same minimum number of missing observations, the most recent window was selected. Years in which no complete 90-day window was available were excluded from further analysis. The selected windows serve as the basis for the annual volatility estimation pipeline.
For each target frequency $f\in{5,10,15}$ minutes, resampling was conducted as follows. The original 1-minute series was first filtered to exclude zero-volume minutes. Aggregated bars were then constructed using standard OHLCV conventions: the Open price corresponds to the first valid price within the interval; the High and Low are, respectively, the maximum and minimum prices observed; the Close is the last valid price; and the Volume is computed as the sum of traded volume over the interval.
To prevent bias from incomplete interval aggregation, only intervals containing exactly $N_{f}$ valid 1-minute observations were retained, where $N_{f}=f$. Intervals with missing observations were discarded.
To ensure high data quality for volatility estimation, a completeness filter was applied after resampling. For each (year, frequency) pair, we computed the completeness ratio as the percentage of expected time intervals for which a valid OHLCV bar was present. Only windows with completeness above 90% were retained:
$\text{Completeness}(Y,f)=\frac{N_{\text{valid}}(Y,f)}{N_{\text{expected}}(f)}\times 100\%$
where $N_{\text{expected}}(f)$ is the theoretical number of bars at frequency $f$ for a 90-day window, and $N_{\text{valid}}(Y,f)$ is the number of valid, fully aggregated bars.
Finally, for each retained window, any remaining observations with missing values were removed via listwise deletion. The resulting datasets consist exclusively of complete, uniformly sampled OHLCV bars. These cleaned series served as direct inputs for volatility estimation and multifractal analysis.
The completeness statistics computed as described above are reported in Table 1.
Table 1: Completeness of selected 90-day Bitcoin data windows (2017-2024).
Year Frequency Valid Bars Expected Bars Completeness (%)
2017 1-min 128,014 129,600 98.78
5-min 24,665 25,920 95.16
10-min 11,883 12,960 91.69
15-min 7,697 8,640 89.09
2018 1-min 129,180 129,600 99.68
5-min 25,543 25,920 98.55
10-min 12,610 12,960 97.30
15-min 8,319 8,640 96.28
2019 1-min 128,999 129,600 99.54
5-min 25,422 25,920 98.08
10-min 12,501 12,960 96.46
15-min 8,205 8,640 94.97
2020 1-min 128,998 129,600 99.54
5-min 25,379 25,920 97.91
10-min 12,461 12,960 96.15
15-min 8,179 8,640 94.66
2021 1-min 129,464 129,600 99.90
5-min 25,795 25,920 99.52
10-min 12,837 12,960 99.05
15-min 8,520 8,640 98.61
2022 1-min 127,584 129,600 98.44
5-min 24,267 25,920 93.62
10-min 11,486 12,960 88.63
15-min 7,280 8,640 84.26
2023 1-min 126,682 129,600 97.75
5-min 23,462 25,920 90.52
10-min 10,784 12,960 83.21
15-min 6,677 8,640 77.28
2024 1-min 128,503 129,600 99.15
5-min 25,052 25,920 96.65
10-min 12,184 12,960 94.01
15-min 7,927 8,640 91.75
A 90% completeness threshold was imposed at each frequency to balance data quality with temporal coverage. For 1-minute and 5-minute data, all years from 2017 to 2024 met this criterion. For 10-minute data, 2022 and 2023 were excluded, while for 15-minute data, 2017, 2022, and 2023 were excluded. The remaining data were sufficient to observe clear patterns, and no contradictory behaviour was seen in partially excluded years at other frequencies. The complete list of retained and excluded years is given in Appendix Table A.1.
No interpolation or imputation was performed at any stage. Missing data were handled exclusively by window selection and exclusion, preserving the integrity of volatility estimates and avoiding artefactual smoothing.
Temporal structure and integrity of each retained window were further assessed. Large temporal gaps, defined as sequences of missing data exceeding one hour, were systematically identified at each frequency. Such gaps were rare for 1-minute and 5-minute data, but more common at coarser resolutions, especially in early years and in 2024. Full gap statistics are provided in Appendix Table A.2.
Integrity checks were conducted to confirm the suitability of the data for volatility modelling. These included confirming the absence of residual missing values after complete-case filtering, ensuring that all closing prices were strictly positive and non-missing, and verifying that timestamp sequences were strictly increasing to preserve temporal alignment.
For each frequency, closing prices are first transformed into log-returns, defined as the first difference of the natural logarithm of consecutive closing prices:
$r_{t}=\log S_{t}-\log S_{t-1}$
where $S_{t}$ denotes the closing price at time $t$. The log-return series is computed separately for each calendar year and each frequency, yielding one series per (year, frequency) pair.
Prior to volatility estimation, each log-return series was screened for data quality issues and extreme values. An absolute threshold test is applied: any return exceeding $|r_{t}|>0.2$ (approximately 22%) is flagged. Statistical outliers are identified using a $z$-score criterion with a threshold of six standard deviations from the mean.
Z-score histograms and Q-Q plots, shown in Appendix Figures B.4-B.6 and B.1-B.3, reveal that Bitcoin log-returns exhibit strong leptokurtosis and significant departures from Gaussianity across all year and frequency configurations. These deviations intensify at higher sampling frequencies, where extreme quantiles diverge markedly from the normal benchmark. The presence of heavy tails suggests that moment-based scaling diagnostics may be sensitive to outliers. To preserve the true structure of return variability, we retain all observations and later apply a shuffle-resampling approach to disentangle heavy-tail effects from temporal dependence in our multifractality tests.
No filtering was applied beyond the initial screening, all flagged returns were retained, and no values were removed or modified.
3.3 Realised Volatility Construction
For each year and frequency configuration, we define a proxy for high-frequency volatility by taking the modulus of log-returns at the original sampling resolution. Let $\{r_{t}\}_{t=1}^{L}$ denote the log-return series over a 90-day window, where $L$ is the number of observations. We define the pointwise volatility proxy as
$RV_{t}=|r_{t}|,\quad t=1,\ldots,L.$
This absolute return series preserves the temporal resolution of the data and captures the local magnitude of return fluctuations. For notational convenience, we refer to $\{RV_{t}\}$ as the realised volatility series, although it is more accurately a proxy for instantaneous volatility rather than an integrated measure.
Formally, for sampling frequency $f$ and year $Y$, let $X(f,Y)=\{r_{t}(f,Y)\}_{t=1}^{L_{f,Y}}$ denote the log-return series. We construct
$RV(f,Y)=\left\{\,|r_{t}(f,Y)|\,\right\}_{t=1}^{L_{f,Y}},$
where $L_{f,Y}$ is the number of valid observations in the selected window. This operation is applied uniformly across all configurations to produce a consistent panel of absolute return series for nonparametric roughness estimation.
We use absolute returns rather than squared returns since the normalised $p$-variation estimator can be directly applied to any $L^{p}$-type measure. The absolute value provides a fat-tailed yet finite input and is more robust to extreme values. Accordingly, we estimate the Hölder roughness of the volatility path itself (as captured by absolute returns), rather than of the price or log-price process. Working directly with the raw volatility magnitude series avoids imposing lognormality or continuous-time modelling assumptions. This follows the approach of Gatheral et al., (2018) and Cont and Das, (2024), who focus on volatility time series for roughness estimation.
3.4 Diagnostic Exploration: Roughness vs $K$
Prior to finalising the roughness estimation procedure, we conduct a diagnostic analysis to examine how the estimated roughness $H_{L,K}$ depends on the number of blocks $K$ (which controls the partition size). This step ensures we choose an appropriate $K$ for each case, as the estimate can be sensitive to this choice. The diagnostic analysis evaluates the behaviour of the roughness estimator $\widehat{H}^{L,K}$ as a function of the number of blocks $K$.
For each calendar year and sampling frequency, $\widehat{H}^{L,K}$ is computed over a wide range of $K$ values, starting from $K=50$, with frequency-specific grid steps to ensure robust coverage of the admissible $K$ range given the sample size. The estimator follows the normalised $p$-variation methodology, computed using block-summed returns and linear regression of $\log W(L,K,p)$ on $1/p$.
Diagnostic plots of $\widehat{H}^{L,K}$ versus $K$ were generated for each (year, frequency) pair. See Section 4.1 for representative results. Across all configurations, these diagnostic plots consistently reveal substantial instability of $\widehat{H}^{L,K}$ as a function of $K$, underscoring the need for principled block size selection.
In principle, a small $K$ leads to highly variable estimates due to insufficient averaging, while a very large $K$ may induce bias by under-representing the high-frequency components of the process. As such, no single value of $K$ ensures consistency in finite samples, and the behaviour of $\widehat{H}^{L,K}$ as a function of $K$ must be carefully examined. Accordingly, the $\widehat{H}^{L,K}$ plot as a function of $K$ is not expected to be flat or perfectly stable, as changes in $K$ inherently induce fluctuations in the estimate due to the bias-variance trade-off.
To mitigate these issues, Cont and Das, (2024) recommend selecting $K$ such that $K\approx\sqrt{L}$, where $L$ is the number of available observations. This choice ensures a trade-off between bias and variance, grounded in the theoretical convergence properties of the normalised $p$-variation statistic:
$\lim_{K\to\infty}\lim_{L\to\infty}W(L,K,p)\to w(x,p,\pi)$ (7)
and maintains a sufficient block size $n=L/K$ to stabilise the estimation.
Following the recommendation of Cont and Das, (2024), we set $K_{\text{opt}}=\lfloor\sqrt{N}\rfloor$, where $N$ is the number of realised volatility observations for the given year and frequency. The corresponding block size is set to $n_{\text{opt}}=K_{\text{opt}}$, ensuring an equal partition. This results in $L=K_{\text{opt}}\times n_{\text{opt}}$ data points used in the estimation. When necessary, a few observations are discarded to make $L$ exactly divisible by $K_{\text{opt}}$. Table 2 reports the resulting partition parameters for each year and frequency.
Table 2: Summary of sample sizes and partitioning parameters $(K,n,L)$ per year and frequency. For each (year, frequency) pair, $K_{\mathrm{opt}}=n_{\mathrm{opt}}=\lfloor\sqrt{N}\rfloor$ and $L=K\times n$ is the number of data points used in estimation.
Frequency Year $N_{\text{points}}$ $K_{\text{opt}}$ $n_{\text{opt}}$ $L$
1-min 2017 128013 357 357 127449
1-min 2018 129179 359 359 128881
1-min 2019 128998 359 359 128881
1-min 2020 128997 359 359 128881
1-min 2021 129463 359 359 128881
1-min 2022 127583 357 357 127449
1-min 2023 126681 355 355 126025
1-min 2024 128502 358 358 128164
5-min 2017 24664 157 157 24649
5-min 2018 25542 159 159 25281
5-min 2019 25421 159 159 25281
5-min 2020 25378 159 159 25281
5-min 2021 25794 160 160 25600
5-min 2022 24266 155 155 24025
5-min 2023 23461 153 153 23409
5-min 2024 25051 158 158 24964
10-min 2017 11882 109 109 11881
10-min 2018 12609 112 112 12544
10-min 2019 12500 111 111 12321
10-min 2020 12460 111 111 12321
10-min 2021 12836 113 113 12769
10-min 2024 12183 110 110 12100
15-min 2018 8318 91 91 8281
15-min 2019 8204 90 90 8100
15-min 2020 8178 90 90 8100
15-min 2021 8519 92 92 8464
15-min 2024 7926 89 89 7921
At the 1-minute frequency, $K_{\text{opt}}$ typically ranges from 355 to 359, whereas at the 15-minute frequency it ranges from 89 to 92. The corresponding number of observations $L$ used in estimation is slightly below the total $N$ due to truncation, which ensures exact partitioning. For example, in 2017 at the 1-minute frequency, only 0.4% of the data is discarded to ensure that $L=K\times n$ holds exactly. These adjustments are minimal and do not materially affect the estimation procedure.
This approach ensures that the roughness estimates rely on stable and theoretically grounded choices of $K$, enabling consistent comparisons across years and frequencies. Given the high-frequency sampling and large sample sizes, the normalised $p$-variation estimator operates within the asymptotic regime necessary for its consistency, as established in Cont and Das, (2024). To assess estimator behaviour, we additionally inspected the curvature and zero-crossing properties of the $\log W(L,K,p)$ curves across a grid of $p$ values. In Section 3.5, we proceed to compute and analyse the final roughness estimates using these $K_{\text{opt}}$ values.
3.5 Final Estimation Setup
[…]
$W(L,K,p)=\frac{1}{K}\sum_{j=1}^{K}|S_{j}|^{p}.$ (8)
To explore the robustness of our estimation and the stability of roughness behaviour, we compute $W(L,K,p)$ over two grids of $p$. The first is a standard grid, defined as $p\in[0.1,4.0]$, as originally proposed by Cont and Das, (2024). The second is an extended exploratory grid (hereafter referred to as the “wide” grid), defined as $p\in[0.01,4.0]$, allowing finer resolution of estimator behaviour at small $p$ values.
For each configuration, we plot $\log W(L,K,p)$ against $1/p$, following standard diagnostic practice. The expected behaviour is a zero-crossing of $\log W(L,K,p)$ at $p=\hat{p}$, which corresponds to the estimated variation index and yields the roughness estimator $\hat{H}=1/\hat{p}$. This diagnostic behaviour is well-defined under the assumption that the process under study admits homogeneous scaling in the sense of Hölder-type regularity. The presence of a clear zero-crossing is a necessary condition for Bitcoin’s volatility to follow a rough volatility scaling law (Hypothesis H1).
Following Cont and Das, (2024), we apply the estimator directly to the series of realised volatility magnitudes, defined as $RV_{t}=|r_{t}|$, where $r_{t}$ is the log-return at time $t$. This approach avoids any transformation to log-volatility or its increments, and the resulting roughness index $\hat{H}$ captures the fine-scale regularity of the volatility path itself. Summary statistics of $\log W(L,K,p)$ behaviour are reported in Table 3 in Section 4.1.
3.6 Diagnostic Tests
3.6.1 Non-Stationarity Diagnostics
To assess whether the failure of the $p$-variation estimator may stem from non-stationarity or structural breaks, we apply three complementary diagnostics to each realised volatility series: the Augmented Dickey-Fuller (ADF) test for mean stationarity, rolling window analysis to evaluate local stability of mean and variance, and structural break detection to identify discrete regime shifts. Together, these methods provide a comprehensive assessment of both gradual and abrupt departures from stationarity.
Each diagnostic is applied independently to all combinations of year and sampling frequency. This convention is adopted throughout the remainder of this subsection and will not be restated for each method.
Unit root tests were conducted using the Augmented Dickey-Fuller (ADF) procedure. For each series, we estimate
$\Delta RV_{t}=\alpha RV_{t-1}+\sum_{i=1}^{p}\beta_{i}\Delta RV_{t-i}+\varepsilon_{t}$
with lag order $p$ selected via the AIC criterion. The resulting test statistics and corresponding critical values are reported in Appendix Table A.3.
Rolling mean and variance stability were assessed by computing the standard deviation of rolling mean and variance estimates. We used a rolling window of length equal to 5% of the sample size (with a minimum of 10 observations). Specifically, we computed
$\operatorname{std}\big{(}\operatorname{mean}_{t-w:t}(RV)\big{)},\quad\operatorname{std}\big{(}\operatorname{var}_{t-w:t}(RV)\big{)}$
where $w$ denotes the rolling window length. The computed statistics are reported in Appendix Table A.4.
Structural breaks were detected using binary segmentation with an $L_{2}$-based cost function. We applied the algorithm with a maximum of five breakpoints per series, solving:
$\min_{\tau_{1},\ldots,\tau_{m}}\sum_{k=0}^{m}\sum_{t=\tau_{k}+1}^{\tau_{k+1}}\left(RV_{t}-\mu_{k}\right)^{2}$
where $\{\tau_{k}\}$ are the breakpoints and $\mu_{k}$ the segment means. The detected breakpoint indices are reported in Appendix Table A.5.
Overall, the applied diagnostics and their corresponding results are documented in Appendix Tables A.3, A.4, and A.5. These results will be interpreted in Section 4.2 to assess Hypothesis H3.
3.6.2 Multifractality Diagnostics
To assess whether Bitcoin’s realised volatility exhibits a multifractal temporal structure, we apply three complementary diagnostics: Multifractal Detrended Fluctuation Analysis (MF-DFA), log-log moment scaling, and wavelet leaders analysis. Each method is implemented systematically across all (year, frequency) combinations and applied uniformly to both log-return and realised volatility series. This convention is maintained throughout and not repeated in the descriptions that follow.
To isolate temporal dependence from distributional effects, we employ a shuffle-based robustness procedure: for each series, the time indices are randomly permuted, thereby removing temporal correlations while preserving the marginal distribution. Comparisons between original and shuffled series allow us to attribute multifractal signatures specifically to dynamic rather than distributional structure.
Multifractal Detrended Fluctuation Analysis (MF-DFA) estimates the generalised Hurst exponent $H(q)$ over the moment range $q\in[-4,4]$, following standard MF-DFA methodology. Specifically, for each $q$, the slope of the log-log regression
$\log F_{q}(s)\sim H(q)\log s$
is computed, where $F_{q}(s)$ denotes the $q$th-order fluctuation function at scale $s$. Summary statistics for the resulting $H(q)$ spectra (including minimum, maximum, mean, and spectral width) are reported in Appendix Table A.6 (log-returns) and Table A.9 (realised volatility).
Log-log moment scaling analysis is used to estimate empirical scaling exponents $\zeta_{q}$ by computing, for each moment order $q\in[-4,4]$, the slope of the scaling relation
$\log M_{q}(\tau)\sim\zeta_{q}\log\tau$
is computed, where $M_{q}(\tau)$ denotes the $q$th-order moment of increments at lag $\tau$. This procedure is implemented over a dense grid of lags $\tau$. Complete results are presented in Appendix Table A.7 (log-returns) and Table A.10 (realised volatility).
Wavelet leaders multifractal analysis estimates the scaling relation for each moment order $q\in[-4,4]$, the scaling relation
$\log_{2}M_{q}(2^{j})\sim\zeta_{q}\,j$
is estimated, where $M_{q}(2^{j})$ denotes the $q$th-order moment of wavelet leaders at dyadic scale $2^{j}$. Scaling exponents $\zeta_{q}$ are obtained via linear regression, and corresponding $R^{2}$ values are recorded to evaluate the quality of fit. Wavelet leaders, constructed from the local suprema of wavelet coefficients across scales, offer robustness to non-stationarities and facilitate the isolation of multiscale fluctuations from deterministic trends. Full results are presented in Appendix Table A.8 (log-returns) and Table A.11 (realised volatility).
A non-linear relationship between $H(q)$ or $\zeta_{q}$ and the moment order $q$ is a hallmark of multifractality, indicating scale-dependent dynamics. The reduction in multifractal spectrum widths observed in the shuffled data serves as a robustness check, indicating that the multifractality originates from temporal correlations rather than from the heavy-tailed distribution of volatility increments.
4 Empirical Results
4.1 Main Result: Roughness Estimation behaviour
We now analyse the behaviour of the normalised $p$-variation statistic $\log W(L,K,p)$ across all (year, frequency) pairs of the dataset.
As detailed in Section 3.4, examining the behaviour of $\widehat{H}^{L,K}$ across varying $K$ offers diagnostic insight into scaling stability and guides the selection of a robust block size.
Figure 1: Estimated roughness index $\widehat{H}^{L,K}$ versus number of blocks $K$ for BTC realised volatility in 2021 at 1-minute (top-left), 5-minute (top-right), 10-minute (bottom-left), and 15-minute (bottom-right) frequencies.
As shown in Figure 1, for the 1-minute data, $\widehat{H}^{L,K}$ is clearly positive for small $K$ (up to approximately $K\approx 200$) and then declines sharply, becoming negative for larger $K$. The 5-minute series exhibits a qualitatively similar pattern, with positive values at low $K$ and a progressive decrease. In contrast, for the 10-minute and 15-minute series, $\widehat{H}^{L,K}$ remains strictly negative throughout the entire $K$-range. In practice, we observed that $\widehat{H}_{L,K}$ tends to be inflated when $K$ is small, reflecting high estimator variance due to coarse partitioning. As $K$ increases, the estimator often fails to return a value, since $\log W(L,K,p)$ remains strictly negative and no zero-crossing is detected. This pattern reflects a fundamental bias-variance trade-off: small $K$ amplifies stochastic variability, while large $K$ discards fine-scale information necessary for identifying path roughness. It emphasises the need for principled selection of $K$, as arbitrary choices can produce misleading indications of roughness at small scales. To address this, the final roughness estimates are based on the theoretically justified criterion $K_{\mathrm{opt}}\approx\sqrt{N}$.
Using the optimal partition size $K_{\mathrm{opt}}\approx\sqrt{N}$, as derived in Section 3.4, we compute the normalised $p$-variation statistic $\log W(L,K,p)$ for all (year, frequency) pairs across a wide range of $p$, following the final estimation procedure of Section 3.5. The summary of minimum and maximum values of $\log W(L,K,p)$ is presented in Table 3.
Table 3: Minimum and maximum values of $\log W(L,K,p)$ for each (frequency, year) pair, for both standard and wide grids.
Frequency Year Standard Grid wide Grid
Min Max Min Max
1min 2017 -1.797 -0.090 -1.797 -0.009
1min 2018 -3.454 -0.110 -3.454 -0.011
1min 2019 -3.149 -0.114 -3.149 -0.011
1min 2020 -1.857 -0.124 -1.857 -0.012
1min 2021 -2.377 -0.098 -2.377 -0.010
1min 2022 -5.021 -0.154 -5.021 -0.015
1min 2023 -6.177 -0.199 -6.177 -0.020
1min 2024 -5.249 -0.175 -5.249 -0.018
5min 2017 -1.636 -0.091 -1.636 -0.009
5min 2018 -3.778 -0.119 -3.778 -0.012
5min 2019 -3.206 -0.120 -3.206 -0.012
5min 2020 -1.911 -0.142 -1.911 -0.014
5min 2021 -2.438 -0.100 -2.438 -0.010
5min 2022 -5.230 -0.154 -5.230 -0.015
5min 2023 -6.375 -0.190 -6.375 -0.019
5min 2024 -5.502 -0.172 -5.502 -0.017
10min 2017 -1.612 -0.090 -1.613 -0.009
10min 2018 -3.866 -0.121 -3.866 -0.012
10min 2019 -3.382 -0.122 -3.382 -0.012
10min 2020 -2.012 -0.147 -2.012 -0.015
10min 2021 -2.577 -0.100 -2.577 -0.010
10min 2024 -5.560 -0.170 -5.560 -0.017
15min 2018 -3.867 -0.123 -3.867 -0.012
15min 2019 -3.358 -0.123 -3.358 -0.012
15min 2020 -2.091 -0.149 -2.091 -0.015
15min 2021 -2.545 -0.100 -2.545 -0.010
15min 2024 -5.544 -0.170 -5.544 -0.017
As shown in Table 3, $\log W(L,K,p)$ remains strictly negative across all years, frequencies, and both the standard and extended $p$-grids. No zero-crossing is observed in any configuration, with maximum values not exceeding $-0.009$, even under the wide grid. This behaviour deviates from the expected pattern under homogeneous roughness, where a unique zero-crossing at $p=\hat{p}$ would yield a well-defined roughness exponent $H=1/\hat{p}$. The consistent absence of such a crossing indicates a systematic failure of the normalised $p$-variation estimator in the Bitcoin setting. The robustness of this result across years and frequencies suggests it reflects a structural property rather than a data artefact.
4.2 Stationarity and Window Stability
We begin by assessing whether non-stationarity could account for the anomalous behaviour of $\log W(L,K,p)$. Unit root tests were performed using the Augmented Dickey-Fuller (ADF) procedure as described in Section 3.6.1. The corresponding test statistics and $p$-values are reported in Appendix Table A.3. In all cases, the null hypothesis of a unit root is strongly rejected at the 5% significance level.
We next assess the local stability of the series by examining the rolling mean and variance statistics, following the procedure outlined in Section 3.6.1. Standard deviations of rolling means and variances were computed for each series, with results summarised in Appendix Table A.4. Across all frequencies and years, the rolling statistics exhibit low variability, indicating stable local behaviour. As expected, dispersion is marginally higher at coarser frequencies (10-minute and 15-minute), reflecting the aggregation of volatility rather than true non-stationarity. Crucially, no systematic trends or regime shifts are detected in the rolling statistics across all years and frequencies.
Finally, we detect structural breaks using binary segmentation with an $L_{2}$-based cost function, as described in Section 3.6.1. The identified breakpoints are reported in Appendix Table A.5. Breakpoints are detected in all series but predominantly align with periods of local volatility clustering. These breakpoints do not correspond to major regime shifts or persistent changes in volatility dynamics. Their occurrence is consistent with stationary behaviour exhibiting local clustering and does not indicate that structural instability accounts for the observed failure of the $p$-variation estimator.
Diagnostic results do not support non-stationarity or structural breaks as plausible explanations for the failure of the normalised $p$-variation estimator. Bitcoin realised volatility appears stationary with stable local dynamics. If non-stationarity were the cause, one would expect to observe systematic trends, level shifts, or persistent regime changes, none of which are observed in the diagnostic results.
4.3 Evidence of Multifractality
The results of the Multifractal Detrended Fluctuation Analysis (MF-DFA), described in Section 3.6.2, are reported in Appendix Tables A.6 and A.9. Across all original series, the estimated $H(q)$ spectra exhibit pronounced non-linearity with respect to $q$, indicating scale-dependent behaviour characteristic of multifractality. For realised volatility, $H(q)$ remain consistently below 1 across all years and sampling frequencies, typically centred around 0.9 for moderate $q$, with spectral widths between 0.06 and 0.42. Log-returns display lower mean exponents but retain comparable curvature in the scaling function. When applied to shuffled series, removing temporal structure while preserving the marginal distribution, the $H(q)$ spectra compress substantially, with spectral widths falling below 0.15 in most cases. This contrast supports the interpretation that the observed multifractality arises from temporal dependence rather than distributional effects.
We next apply the log-log moment scaling analysis, an alternative multifractality diagnostic based on the scaling behaviour of $\zeta_{q}$, as described in Section 3.6.2. Estimated $\zeta_{q}$ spectra for realised volatility and log-returns are presented in Appendix Tables A.7 and A.10. For realised volatility, the $\zeta_{q}$ spectra exhibit pronounced curvature in $q$, with widths ranging from 0.04 to over 2.0 and mean $R^{2}$ values generally exceeding 0.85, indicating stable scaling across temporal resolutions. Log-return series show more heterogeneity (especially at lower sampling frequencies) but maintain clear non-linearity in the $\zeta_{q}$ profile for moderate to large $q$. In the shuffled series, the $\zeta_{q}$ spectra flatten substantially, with widths typically below 0.2 and minimal curvature. These patterns are consistent with a temporal origin of the observed multifractality in both series.
The wavelet leaders multifractal analysis offers a robust characterisation of scaling behaviour, particularly in the presence of trends or non-stationarities. Estimated $\zeta_{q}$ exponents for both realised volatility and log-return series are presented in Appendix Tables A.8 and A.11. For realised volatility, the $\zeta_{q}$ spectra exhibit consistent curvature with spectral widths between 0.04 and 0.51 and mean $R^{2}$ values typically exceeding 0.85. Log-return series show similar but generally weaker curvature, with greater sensitivity to sampling resolution. In the shuffled series, spectral widths decline substantially (often by more than 50%) and curvature is notably reduced.
The MF-DFA, log-log moment scaling, and wavelet leaders analyses consistently indicate that Bitcoin realised volatility exhibits a multifractal temporal structure. This conclusion is supported by the marked reduction in spectral curvature and width observed in shuffled controls, which preserve the marginal distribution but remove temporal dependence. These results suggest that the observed multifractality arises from temporal correlations rather than from heavy-tailed marginal distributions. This violates the homogeneity assumptions required by the normalised $p$-variation framework and provides a structural explanation for the failure to identify a well-defined scaling exponent. The apparent roughness at small $q$ reflects localised scaling properties embedded in a broader multifractal structure, rather than a stable global Hurst exponent.
4.4 Hypothesis Evaluation
H1. Bitcoin volatility does not admit a global scaling law compatible with the normalised $p$-variation framework:
Our primary empirical result, the strict negativity of $\log W(L,K,p)$ across all years, frequencies, and $p$-grids, demonstrates the absence of any stable scaling law compatible with the normalised $p$-variation estimator in Bitcoin realised volatility. This supports the interpretation that the apparent roughness reported in prior literature does not reflect an intrinsic property of the volatility process, but rather the failure of the $p$-variation framework to capture consistent scaling behaviour in this context. These findings confirm Hypothesis H1, as the strict negativity of $\log W(L,K,p)$ across all configurations indicates the absence of a well-defined scaling law.
H2. Estimation Artefacts and Finite-Sample Effects:
While finite-sample effects and estimation artefacts may contribute to the instability of the roughness estimator, they cannot fully account for the systematic and robust failure observed here. The optimal partitioning strategy $K\approx\sqrt{N}$, validated in prior studies Cont and Das, (2024), was applied consistently, and window stability diagnostics confirm that the input series meet the methodological requirements of the estimator. Market microstructure noise, irregular liquidity, and large volatility jumps, which are common in Bitcoin markets, likely exacerbate finite-sample distortions and further undermine the reliability of roughness estimation. However, finite-sample effects alone are insufficient to account for the observed empirical anomaly. We therefore reject $H_{2}$ as a primary explanation, since finite-sample effects and estimation artefacts do not account for the systematic failure of $\log W(L,K,p)$.
H3. Non-Stationarity or Regime Shifts:
Our suite of stationarity diagnostics, including the Augmented Dickey-Fuller test for unit roots, rolling mean and variance statistics for local stability, and structural break detection via binary segmentation with an $L_{2}$-based cost function, yields no evidence that non-stationarity or persistent structural instability accounts for the failure of the $p$-variation estimator. The realised volatility series are stationary over the studied periods, and the detected breakpoints correspond to local volatility clustering rather than structural regime changes. We therefore reject $H_{3}$ as a primary explanation for the observed behaviour of $\log W(L,K,p)$, given the absence of non-stationarity or structural instability.
H4. Multifractality of Bitcoin Volatility:
The multifractality hypothesis is decisively supported by our results. All three diagnostic approaches, MF-DFA, log-log moment scaling, and wavelet leaders analysis, exhibit consistent non-linear scaling in the original series and show substantial collapse of multifractal indicators under shuffle-based controls. This confirms that the observed multifractality in Bitcoin volatility, particularly in realised volatility, arises from genuine temporal correlations rather than from heavy-tailed marginal distributions. The presence of this multifractal temporal structure implies non-linear, scale-dependent behaviour that directly violates the homogeneity assumption of the normalised $p$-variation framework. We therefore accept $H_{4}$ as the most plausible explanation for the observed behaviour of $\log W(L,K,p)$, as the multifractal structure of Bitcoin volatility is supported by all diagnostic results and offers a coherent account of the estimator’s failure.
5 Discussion
5.1 Structural Interpretation
Our analysis demonstrates that Bitcoin realised volatility exhibits pronounced scale dependence and temporal heterogeneity, in contrast to the patterns typically observed in traditional financial assets. Evidence from multiple diagnostics, reinforced by shuffled controls, consistently indicates the presence of multifractality and the absence of a global scaling regime. This empirical structure is incompatible with the assumption of homogeneous Hölder continuity underlying the rough volatility framework.
The multifractality observed in Bitcoin likely arises from structural features specific to cryptocurrency markets. These include fragmented liquidity, abrupt price movements, and regime shifts shaped by technological and regulatory developments. Such conditions generate heterogeneous trading behaviour and fluctuating activity levels, which are established drivers of multifractal scaling in financial time series Bacry et al., (2001); Calvet and Fisher, (2002). As a result, Bitcoin volatility displays local scaling properties that vary over time and across frequencies, precluding the use of a single global roughness exponent.
This structural heterogeneity explains the empirical failure of roughness estimation techniques in our study. Rather than indicating a methodological flaw, the estimator’s breakdown underscores the fundamental incompatibility between rough volatility assumptions and the multifractal structure of Bitcoin markets.
5.2 Implications for Modelling
The structural properties of Bitcoin volatility identified in this study raise concerns about the theoretical validity of applying rough volatility models in this setting. These models rely on the presence of a global scaling regime and homogeneous Hölder continuity. However, empirical findings show that Bitcoin volatility exhibits scale-dependent dynamics and temporal heterogeneity, which violate these fundamental assumptions. Consequently, the existence of a well-defined roughness index cannot be substantiated for Bitcoin.
This incompatibility motivates a broader reconsideration of modelling frameworks for assets that exhibit multifractal characteristics. Models that enforce constant scaling behaviour or a single global regularity parameter are not suitable for processes governed by local, time-varying dynamics. In particular, estimation techniques such as the normalised $p$-variation Cont and Das, (2024) and log-log regressions Fukasawa et al., (2022); Rogers, (2019) lose both interpretability and statistical consistency when applied to processes lacking global scaling symmetry.
In contrast, multifractal modelling frameworks offer a more theoretically coherent basis for capturing the empirical behaviour of Bitcoin volatility Calvet and Fisher, (2002); Bacry et al., (2001). Approaches such as the multifractal random walk Bacry et al., (2001) and multifractal stochastic volatility Calvet and Fisher, (2002) allow for non-linear scaling and dynamic heterogeneity without requiring a constant roughness parameter. These models are better aligned with the mathematical properties observed in the data and avoid the structural misspecification inherent in rough volatility assumptions Cont and Das, (2024).
5.3 Practical Consequences
The failure of normalised $p$-variation estimation, coupled with the presence of multifractality in Bitcoin volatility, has significant implications for practical applications in forecasting, pricing, and risk management. These findings demonstrate that methods relying on global scaling assumptions are unreliable in the context of cryptocurrency markets.
In practical applications, volatility models used for calibration and prediction must be validated within the specific asset class. In the case of Bitcoin, the absence of a stable roughness index leads to unstable parameter estimates and poor out-of-sample performance when rough volatility models are employed. Volatility models calibrated to equity or foreign exchange markets Gatheral et al., (2018) are not directly transferable to cryptocurrencies, given the structural differences in volatility dynamics.
Reliable volatility forecasting and option pricing for Bitcoin require models that can capture its scale-dependent and time-varying structure. Standard approaches, including GARCH-type models and those assuming constant Hölder regularity, tend to underestimate long memory and tail risk. In contrast, multifractal stochastic volatility models Calvet and Fisher, (2002) and wavelet-based methods offer more robust predictive performance by capturing the non-linear and scale-dependent dependencies observed in Bitcoin data.
Practical implementation must also account for market microstructure effects. Bitcoin markets exhibit fragmented liquidity, irregular order flow, and evolving conditions that induce local scaling irregularities and long-range dependence. These features are not captured by standard continuous-path diffusion models. Effective applications such as market-making, hedging, and intraday risk management require models that incorporate structural elements including jumps, regime shifts, and fluctuations in trading activity. Neglecting these factors leads to distorted risk assessments and degraded model performance in high-frequency or volatile environments.
5.4 Limitations and Future Work
While this study highlights the limitations of rough volatility estimation for Bitcoin and confirms the presence of multifractal structure, several limitations should be acknowledged. The study focuses exclusively on realised volatility derived from high-frequency price data. Although well-suited for $p$-variation analysis, this measure captures only a subset of the latent volatility process. The scaling behaviour of alternative measures, such as implied volatility or state-space estimates, has not been examined. Extending the analysis to these measures would clarify whether multifractality is an intrinsic property of Bitcoin volatility or an artefact of realised volatility estimation, and would offer insights into how market expectations reflect multiscale behaviour.
Second, despite careful preprocessing and partitioning following best practices Cont and Das, (2024), our results remain subject to finite-sample limitations. While stationarity and stability diagnostics mitigate concerns about artefactual distortions, the evolving and volatile nature of Bitcoin’s microstructure may still introduce residual biases. Future research could assess the robustness of these findings under alternative sampling schemes, adaptive windowing methods, or extended time horizons (60-day or 120-day windows) to evaluate sensitivity to temporal resolution and aggregation scale.
Third, while our multifractality diagnostics, MF-DFA, log-log scaling, and wavelet leaders, yield consistent results, they are descriptive and empirical. Rigorously confirming multifractal dynamics and distinguishing them from noise or artefacts requires formal statistical testing, including confidence intervals and significance assessments. Approaches such as Wendt et al., (2007) provide a promising foundation. Applying these methods to high-frequency crypto data would constitute a valuable methodological contribution, strengthening the empirical validity of findings in this domain.
These results motivate the development of multifractal or multiscale volatility models tailored to the structural features of Bitcoin markets. Extending the analysis to other crypto-assets such as Ethereum and to traditional markets such as foreign exchange and equities would help determine whether the observed scaling behaviour is specific to Bitcoin or reflects broader properties common to digital assets. Such comparisons would contribute to a more unified understanding of volatility across diverse markets.
6 Conclusion
This study assessed the applicability of rough volatility methodologies to Bitcoin using the normalised $p$-variation framework proposed by Cont and Das, (2024). The statistic $\log W(L,K,p)$ remains strictly negative across all configurations, precluding the estimation of a valid roughness index. Diagnostic tests rule out non-stationarity, structural breaks, and sampling artefacts as plausible explanations. Instead, consistent evidence from MF-DFA, log-log moment scaling, and wavelet leaders indicates that Bitcoin realised volatility follows a robust multifractal temporal structure, driven by dynamic correlations rather than marginal distributional effects. This violates the homogeneity assumptions required by rough volatility models and accounts for the systematic failure of the estimator.
These findings demonstrate that Bitcoin volatility cannot be adequately described by a single roughness parameter or Hurst exponent. This result is robust across all frequencies, time periods, and estimation settings. Consequently, rough volatility models are structurally incompatible with Bitcoin and are likely to produce unreliable forecasts and pricing outcomes. Models that incorporate multifractal or scale-dependent behaviour, such as multifractal stochastic volatility models, offer a more appropriate empirical fit. Future research should extend this analysis to alternative volatility measures, asset classes, and modelling frameworks to foster a broader understanding of volatility across heterogeneous markets.
Appendix A: Supplementary Tables
Table A.1: Retained and excluded years by frequency after applying 90% completeness threshold (2017–2024).
Frequency Years Retained
1-min 2017, 2018, 2019, 2020, 2021, 2022, 2023, 2024
5-min 2017, 2018, 2019, 2020, 2021, 2022, 2023, 2024
10-min 2017, 2018, 2019, 2020, 2021, 2024
15-min 2018, 2019, 2020, 2021, 2024
Frequency Years Excluded (Below 90%)
1-min None
5-min None
10-min 2022, 2023
15-min 2017, 2022, 2023
Table A.2: Summary of missing data and large gaps ($>$1 hour) by frequency and year in retained 90-day BTC windows.
Frequency Year Total Missing Minutes Gaps $>$1 Hour (Count)
1-min 2017 1586 0
1-min 2018 420 0
1-min 2019 601 1
1-min 2020 602 0
1-min 2021 136 0
1-min 2022 2016 1
1-min 2023 2918 1
1-min 2024 1097 2
5-min 2017 6275 0
5-min 2018 1885 0
5-min 2019 2490 1
5-min 2020 2705 0
5-min 2021 625 0
5-min 2022 8265 3
5-min 2023 12285 3
5-min 2024 4340 2
10-min 2017 10770 10
10-min 2018 3500 0
10-min 2019 4590 1
10-min 2020 4990 3
10-min 2021 1230 0
10-min 2024 7760 4
15-min 2018 4815 0
15-min 2019 6525 3
15-min 2020 6915 6
15-min 2021 1800 0
15-min 2024 10695 13
Table A.3: ADF Stationarity Summary Table
Frequency Year $N_{\text{obs}}$ ADF_stat $p$-value Stationary
1min 2017 128013 -17.268377 $5.88\times 10^{-30}$ True
1min 2018 129179 -21.103397 $0.00$ True
1min 2019 128998 -22.996778 $0.00$ True
1min 2020 128997 -16.217826 $3.93\times 10^{-29}$ True
1min 2021 129463 -19.473206 $0.00$ True
1min 2022 127583 -20.992987 $0.00$ True
1min 2023 126681 -20.199538 $0.00$ True
1min 2024 128502 -19.502510 $0.00$ True
5min 2017 24664 -10.745265 $2.75\times 10^{-19}$ True
5min 2018 25542 -12.375319 $5.17\times 10^{-23}$ True
5min 2019 25421 -12.273721 $8.53\times 10^{-23}$ True
5min 2020 25378 -10.130513 $8.94\times 10^{-18}$ True
5min 2021 25794 -10.793006 $2.11\times 10^{-19}$ True
5min 2022 24266 -13.884217 $6.13\times 10^{-26}$ True
5min 2023 23461 -17.513342 $4.32\times 10^{-30}$ True
5min 2024 25051 -11.923340 $4.97\times 10^{-22}$ True
10min 2017 11882 -7.661084 $1.69\times 10^{-11}$ True
10min 2018 12609 -8.932990 $9.70\times 10^{-15}$ True
10min 2019 12500 -9.974204 $2.20\times 10^{-17}$ True
10min 2020 12460 -7.922876 $3.68\times 10^{-12}$ True
10min 2021 12836 -8.282212 $4.49\times 10^{-13}$ True
10min 2024 12183 -12.177698 $1.37\times 10^{-22}$ True
15min 2018 8318 -7.222499 $2.09\times 10^{-10}$ True
15min 2019 8204 -8.225838 $6.25\times 10^{-13}$ True
15min 2020 8178 -7.227016 $2.04\times 10^{-10}$ True
15min 2021 8519 -7.315916 $1.23\times 10^{-10}$ True
15min 2024 7926 -8.943357 $9.13\times 10^{-15}$ True
Table A.4: Rolling Mean/Variance Stability Table
(Standard deviation of rolling mean/variance; lower = more stable)
Frequency Year $N_{\text{obs}}$ RollingWindow RollingMeanStd RollingVarStd
1min 2017 128013 6400 0.000487 $1.69\times 10^{-6}$
1min 2018 129179 6458 0.000276 $5.53\times 10^{-7}$
1min 2019 128998 6449 0.000310 $9.69\times 10^{-7}$
1min 2020 128997 6449 0.000537 $3.24\times 10^{-6}$
1min 2021 129463 6473 0.000322 $1.00\times 10^{-6}$
1min 2022 127583 6379 0.000148 $2.63\times 10^{-7}$
1min 2023 126681 6334 0.000136 $1.93\times 10^{-7}$
1min 2024 128502 6425 0.000176 $2.61\times 10^{-7}$
5min 2017 24664 1233 0.001242 $1.00\times 10^{-5}$
5min 2018 25542 1277 0.000629 $2.72\times 10^{-6}$
5min 2019 25421 1271 0.000766 $5.94\times 10^{-6}$
5min 2020 25378 1268 0.001294 $1.85\times 10^{-5}$
5min 2021 25794 1289 0.000762 $4.88\times 10^{-6}$
5min 2022 24266 1213 0.000355 $1.48\times 10^{-6}$
5min 2023 23461 1173 0.000315 $1.12\times 10^{-6}$
5min 2024 25051 1252 0.000394 $1.23\times 10^{-6}$
10min 2017 11882 594 0.001835 $2.01\times 10^{-5}$
10min 2018 12609 630 0.000910 $5.43\times 10^{-6}$
10min 2019 12500 625 0.001055 $1.71\times 10^{-5}$
10min 2020 12460 623 0.001861 $3.78\times 10^{-5}$
10min 2021 12836 641 0.001109 $9.78\times 10^{-6}$
10min 2024 12183 609 0.000546 $2.65\times 10^{-6}$
15min 2018 8318 415 0.001161 $8.16\times 10^{-6}$
15min 2019 8204 410 0.001331 $2.25\times 10^{-5}$
15min 2020 8178 408 0.002326 $6.73\times 10^{-5}$
15min 2021 8519 425 0.001435 $1.72\times 10^{-5}$
15min 2024 7926 396 0.000666 $3.50\times 10^{-6}$
Table A.5: Structural Breakpoints Detected by Binary Segmentation (Maximum 5 per Series)
Frequency Year $N_{\text{obs}}$Breaks Break Indices
1min 2017 $128,013$ 5 [81,225, 93,345, 98,420, 110,855, 118,520]
1min 2018 $129,179$ 5 [20,100, 32,830, 50,455, 64,455, 69,100]
1min 2019 $128,998$ 5 [10,335, 60,520, 66,495, 77,840, 100,395]
1min 2020 $128,997$ 5 [100,870, 101,400, 101,440, 103,445, 118,760]
1min 2021 $129,463$ 5 [2,475, 14,290, 18,080, 42,755, 92,420]
1min 2022 $127,583$ 5 [25,405, 34,420, 70,240, 75,450, 103,895]
1min 2023 $126,681$ 5 [5,480, 26,570, 76,555, 85,755, 107,190]
1min 2024 $128,502$ 5 [51,000, 75,290, 87,365, 118,305, 118,780]
5min 2017 $24,664$ 5 [15,335, 17,530, 18,760, 21,230, 22,790]
5min 2018 $25,542$ 5 [3,960, 6,475, 6,495, 9,870, 14,860]
5min 2019 $25,421$ 5 [2,060, 11,985, 13,180, 15,440, 19,885]
5min 2020 $25,378$ 5 [19,980, 20,085, 20,090, 20,535, 23,675]
5min 2021 $25,794$ 5 [495, 2,855, 3,605, 8,195, 8,490]
5min 2022 $24,266$ 5 [4,760, 6,785, 13,240, 14,580, 19,765]
5min 2023 $23,461$ 5 [995, 4,715, 15,780, 18,055, 19,885]
5min 2024 $25,051$ 5 [9,810, 14,640, 17,065, 23,055, 23,145]
10min 2017 $11,882$ 5 [7,220, 7,495, 10,165, 10,480, 10,945]
10min 2018 $12,609$ 5 [2,060, 3,185, 3,195, 4,890, 7,305]
10min 2019 $12,500$ 5 [890, 900, 1,030, 5,925, 9,840]
10min 2020 $12,460$ 5 [9,875, 10,000, 10,035, 10,155, 11,645]
10min 2021 $12,836$ 5 [245, 1,430, 1,600, 1,830, 4,240]
10min 2024 $12,183$ 5 [4,700, 7,105, 8,300, 11,205, 11,255]
15min 2018 $8,318$ 5 [1,305, 2,095, 2,105, 3,185, 4,800]
15min 2019 $8,204$ 5 [680, 3,905, 4,305, 4,965, 6,505]
15min 2020 $8,178$ 5 [6,550, 6,600, 6,620, 6,675, 7,565]
15min 2021 $8,519$ 5 [160, 950, 1,065, 1,205, 2,825]
15min 2024 $7,926$ 5 [3,010, 3,905, 4,610, 5,380, 7,195]
Table A.6: Summary of MF-DFA Scaling Exponents $H(q)$ for Bitcoin Log-Returns (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) of $H(q)$ spectra.
Freq. Year Original $H(q)$ Shuffled $H(q)$
Min Max Mean Width Min Max Mean Width
1min 2017 0.404 2.167 1.148 1.763 0.468 0.530 0.501 0.061
1min 2018 0.442 0.610 0.528 0.168 0.472 0.540 0.509 0.068
1min 2019 0.356 0.595 0.515 0.239 0.417 0.539 0.504 0.122
1min 2020 0.368 0.572 0.517 0.204 0.418 0.540 0.507 0.122
1min 2021 0.410 0.556 0.503 0.146 0.470 0.540 0.510 0.070
1min 2022 0.443 0.587 0.524 0.144 0.460 0.541 0.508 0.081
1min 2023 0.448 0.562 0.515 0.114 0.463 0.570 0.515 0.107
1min 2024 0.419 0.616 0.523 0.197 0.444 0.521 0.490 0.077
5min 2017 0.392 0.614 0.509 0.222 0.502 0.593 0.540 0.091
5min 2018 0.425 0.608 0.527 0.183 0.480 0.546 0.518 0.066
5min 2019 0.345 0.617 0.511 0.272 0.463 0.586 0.529 0.123
5min 2020 0.332 0.592 0.498 0.260 0.439 0.635 0.547 0.196
5min 2021 0.405 0.571 0.510 0.166 0.463 0.557 0.515 0.094
5min 2022 0.453 0.585 0.527 0.132 0.458 0.549 0.515 0.091
5min 2023 0.455 0.730 0.555 0.275 0.436 0.563 0.499 0.127
5min 2024 0.412 0.642 0.525 0.230 0.453 0.549 0.507 0.096
10min 2017 0.379 0.617 0.522 0.237 0.526 0.610 0.564 0.084
10min 2018 0.422 0.622 0.533 0.200 0.497 0.593 0.545 0.096
10min 2019 0.339 0.635 0.520 0.296 0.397 0.594 0.526 0.197
10min 2020 0.365 0.611 0.512 0.246 0.402 0.636 0.539 0.234
10min 2021 0.403 0.579 0.515 0.176 0.446 0.530 0.499 0.084
10min 2024 0.400 0.662 0.527 0.262 0.473 0.592 0.537 0.119
15min 2018 0.429 0.637 0.533 0.208 0.506 0.601 0.552 0.095
15min 2019 0.365 0.645 0.523 0.281 0.448 0.625 0.545 0.177
15min 2020 0.342 0.630 0.506 0.288 0.370 0.653 0.540 0.283
15min 2021 0.394 0.586 0.511 0.192 0.466 0.555 0.513 0.089
15min 2024 0.402 0.653 0.527 0.251 0.452 0.564 0.512 0.112
Table A.7: Summary of Log-Log Scaling Exponents for Bitcoin Log-Returns (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) of scaling coefficients.
Freq. Year Original Shuffled
Min Max Mean Width Min Max Mean Width
1min 2017 -4.919 0.005 -0.963 4.924 0.000 0.337 0.102 0.337
1min 2018 -3.914 0.003 -0.796 3.917 0.000 0.500 0.150 0.500
1min 2019 -2.798 0.004 -0.564 2.802 0.000 0.343 0.086 0.343
1min 2020 -2.964 0.007 -0.555 2.971 -0.001 0.554 0.108 0.555
1min 2021 -0.064 1.407 0.229 1.471 -0.036 $0$0.108 0.036
1min 2022 -2.379 0.005 -0.654 2.384 0.000 0.593 0.117 0.593
1min 2023 -0.034 0.009 -0.235 0.043 0.000 0.177 0.042 0.177
1min 2024 -1.420 0.007 -0.408 1.427 -0.001 0.083 0.016 0.083
5min 2017 -1.812 0.007 -0.494 1.819 -0.001 0.959 0.235 0.960
5min 2018 -0.014 0.003 -0.123 0.017 -0.005 0.242 0.047 0.247
5min 2019 -0.105 0.533 0.130 0.638 -0.002 $0$0.073 0.002
5min 2020 -2.053 0.009 -0.487 2.062 -0.000 1.004 0.293 1.004
5min 2021 -0.058 0.007 -0.103 0.065 -0.002 0.065 0.021 0.067
5min 2022 -0.023 0.005 -0.106 0.028 0.000 0.999 0.094 0.999
5min 2023 -0.021 0.009 -0.338 0.030 -0.001 0.010 0.003 0.010
5min 2024 -0.041 0.007 -0.406 0.048 -0.010 -0.312 -0.063 0.302
10min 2017 -0.084 0.068 0.008 0.152 -0.949 $0$0.333 0.949
10min 2018 -0.016 0.135 0.029 0.151 -0.115 $0$0.045 0.115
10min 2019 -0.360 0.002 -0.079 0.362 -0.168 0.001 -0.048 0.168
10min 2020 -1.166 0.008 -0.333 1.174 -0.345 0.007 -0.078 0.352
10min 2021 -1.341 0.004 -0.303 1.345 -0.688 0.005 -0.102 0.693
10min 2024 -1.239 0.006 -0.304 1.245 -0.007 0.197 0.031 0.204
15min 2018 -1.732 0.001 -0.391 1.733 -0.205 0.004 -0.036 0.209
15min 2019 -0.019 0.001 -0.107 0.021 -0.633 0.029 -0.129 0.662
15min 2020 -0.127 1.682 0.284 1.809 0.000 0.495 0.151 0.495
15min 2021 -0.087 0.006 -0.119 0.093 -0.009 0.124 0.029 0.133
15min 2024 -0.062 0.006 -0.225 0.068 -0.009 0.262 0.035 0.271
Table A.8: Summary of Wavelet Leaders Scaling Exponents for Bitcoin Log-Returns (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) for each case.
Freq. Year Original Shuffled
Min Max Mean Width Min Max Mean Width
1min 2017 -0.200 0.063 -0.036 0.263 -0.167 -0.001 -0.119 0.166
1min 2018 -0.190 0.060 -0.024 0.250 -0.146 -0.054 -0.110 0.092
1min 2019 -0.245 0.047 -0.078 0.292 -0.152 -0.069 -0.098 0.083
1min 2020 -0.249 0.124 -0.034 0.373 -0.144 -0.046 -0.096 0.098
1min 2021 -0.199 0.140 -0.096 0.339 -0.122 -0.008 -0.056 0.114
1min 2022 -0.195 -0.015 -0.094 0.180 -0.094 -0.003 -0.037 0.091
1min 2023 -0.123 0.146 0.017 0.269 -0.157 -0.071 -0.110 0.086
1min 2024 -0.199 0.043 -0.057 0.242 -0.164 -0.050 -0.094 0.114
5min 2017 -0.217 0.079 -0.039 0.297 -0.218 -0.008 -0.154 0.210
5min 2018 -0.220 0.002 -0.069 0.222 -0.044 0.016 -0.008 0.060
5min 2019 -0.252 0.073 -0.043 0.325 -0.206 -0.103 -0.143 0.103
5min 2020 -0.236 0.137 -0.010 0.372 -0.316 -0.001 -0.116 0.315
5min 2021 -0.181 -0.060 -0.107 0.121 -0.199 -0.082 -0.120 0.117
5min 2022 -0.114 0.039 -0.032 0.153 -0.146 -0.031 -0.086 0.115
5min 2023 -0.118 0.061 0.004 0.179 -0.196 -0.149 -0.159 0.047
5min 2024 -0.115 0.105 0.016 0.220 -0.249 -0.007 -0.158 0.242
10min 2017 -0.015 0.189 0.074 0.204 -0.060 0.042 0.000 0.102
10min 2018 -0.215 0.009 -0.088 0.225 -0.185 0.001 -0.067 0.186
10min 2019 -0.143 0.124 -0.018 0.267 -0.174 0.118 -0.020 0.292
10min 2020 -0.039 0.165 0.046 0.204 -0.249 0.127 -0.037 0.377
10min 2021 -0.188 -0.040 -0.104 0.148 -0.114 -0.023 -0.058 0.091
10min 2024 -0.268 0.093 -0.050 0.362 -0.019 0.185 0.061 0.204
15min 2018 -0.194 -0.021 -0.082 0.173 -0.008 0.113 0.040 0.121
15min 2019 -0.228 0.059 -0.066 0.288 -0.383 -0.037 -0.176 0.346
15min 2020 -0.037 0.190 0.049 0.227 -0.160 0.130 -0.012 0.290
15min 2021 -0.159 -0.067 -0.109 0.092 -0.199 -0.003 -0.087 0.197
15min 2024 -0.220 0.081 -0.037 0.301 -0.009 0.041 0.018 0.050
Table A.9: Summary of MF-DFA Scaling Exponents $H(q)$ for Bitcoin Realised Volatility (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) of $H(q)$ spectra.
Freq. Year Original $H(q)$ Shuffled $H(q)$
Min Max Mean Width Min Max Mean Width
1min 2017 0.863 2.082 1.351 1.219 0.460 0.550 0.504 0.090
1min 2018 0.812 0.867 0.848 0.055 0.487 0.572 0.531 0.085
1min 2019 0.748 0.912 0.876 0.163 0.395 0.558 0.510 0.163
1min 2020 0.871 0.945 0.896 0.074 0.422 0.618 0.534 0.195
1min 2021 0.829 0.920 0.894 0.091 0.457 0.559 0.517 0.102
1min 2022 0.781 0.939 0.872 0.158 0.453 0.579 0.526 0.126
1min 2023 0.780 0.931 0.867 0.151 0.438 0.581 0.518 0.143
1min 2024 0.814 1.023 0.921 0.209 0.462 0.572 0.527 0.110
5min 2017 0.850 0.918 0.901 0.068 0.470 0.591 0.533 0.121
5min 2018 0.779 0.904 0.854 0.125 0.460 0.594 0.530 0.134
5min 2019 0.642 0.948 0.837 0.306 0.411 0.612 0.527 0.200
5min 2020 0.814 0.942 0.885 0.128 0.425 0.676 0.549 0.251
5min 2021 0.820 0.907 0.886 0.087 0.439 0.578 0.517 0.139
5min 2022 0.726 0.918 0.822 0.192 0.410 0.582 0.526 0.172
5min 2023 0.730 1.040 0.871 0.310 0.428 0.595 0.523 0.167
5min 2024 0.791 0.969 0.876 0.178 0.393 0.563 0.508 0.170
10min 2017 0.840 0.946 0.914 0.106 0.494 0.574 0.535 0.080
10min 2018 0.726 0.889 0.834 0.163 0.457 0.593 0.527 0.135
10min 2019 0.592 0.917 0.835 0.324 0.321 0.627 0.526 0.306
10min 2020 0.824 0.863 0.870 0.040 0.352 0.717 0.536 0.365
10min 2021 0.777 0.882 0.862 0.104 0.440 0.586 0.526 0.146
10min 2024 0.767 0.951 0.880 0.184 0.419 0.607 0.516 0.188
15min 2018 0.755 0.912 0.837 0.157 0.475 0.632 0.528 0.157
15min 2019 0.646 0.948 0.834 0.302 0.383 0.667 0.523 0.284
15min 2020 0.758 0.840 0.820 0.082 0.322 0.744 0.527 0.422
15min 2021 0.755 0.874 0.842 0.119 0.419 0.612 0.527 0.193
15min 2024 0.747 0.942 0.870 0.195 0.478 0.654 0.527 0.176
Table A.10: Summary of Log-Log Scaling Exponents for Bitcoin Realised Volatility (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) of scaling coefficients.
Freq. Year Original Shuffled
Min Max Mean Width Min Max Mean Width
1min 2017 -5.001 0.066 -0.927 5.067 -0.001 0.338 0.094 0.339
1min 2018 -4.314 0.049 -0.849 4.363 -0.002 1.272 0.328 1.274
1min 2019 -3.714 0.082 -0.754 3.796 -0.002 $0$0.096 0.002
1min 2020 -2.628 0.122 -0.451 2.750 -0.000 0.290 0.073 0.290
1min 2021 0.107 1.084 0.446 0.976 -0.008 0.465 0.116 0.473
1min 2022 -2.086 0.054 -0.557 2.140 -0.001 0.569 0.113 0.570
1min 2023 -0.655 0.064 -0.156 0.719 -0.000 0.306 0.055 0.306
1min 2024 -0.933 0.110 -0.208 1.043 -0.001 0.014 0.002 0.015
5min 2017 -3.164 0.101 -0.629 3.265 -0.000 1.204 0.339 1.204
5min 2018 -0.279 0.057 -0.063 0.336 -0.010 -$0$0.071 0.010
5min 2019 -0.014 1.010 0.281 1.025 -0.015 0.125 0.026 0.140
5min 2020 -1.702 0.181 -0.283 1.883 0.000 0.411 0.091 0.411
5min 2021 -0.336 0.110 -0.011 0.446 -0.007 0.176 0.022 0.183
5min 2022 -0.359 0.041 -0.059 0.400 0.000 0.220 0.060 0.220
5min 2023 -1.919 0.066 -0.311 1.985 -0.001 0.367 0.068 0.368
5min 2024 -1.503 0.079 -0.202 1.582 -0.007 -$0$0.130 0.007
10min 2017 -0.006 0.131 0.045 0.137 -0.793 $0$0.195 0.793
10min 2018 -0.121 0.049 -0.022 0.170 -0.561 $0$0.130 0.561
10min 2019 -0.024 0.063 0.018 0.087 0.000 0.500 0.133 0.500
10min 2020 -0.548 0.126 -0.067 0.674 -0.328 0.003 -0.061 0.331
10min 2021 -1.739 0.095 -0.354 1.835 -0.163 0.009 -0.018 0.172
10min 2024 -1.962 0.074 -0.281 2.036 -0.000 0.168 0.042 0.168
15min 2018 -1.505 0.035 -0.305 1.540 -0.002 0.041 0.013 0.043
15min 2019 -0.004 0.590 0.142 0.594 -1.786 0.010 -0.377 1.796
15min 2020 -0.010 0.996 0.317 1.006 0.000 0.236 0.067 0.236
15min 2021 -0.006 0.286 0.094 0.292 0.000 0.145 0.025 0.145
15min 2024 -1.249 0.091 -0.193 1.340 -0.002 0.152 0.038 0.154
Table A.11: Summary of Wavelet Leaders Scaling Exponents for Bitcoin Realised Volatility (Original vs. Shuffled, 2017–2024). Table reports minimum, maximum, mean, and width ($\max-\min$) for each case.
Freq. Year Original Shuffled
Min Max Mean Width Min Max Mean Width
1min 2017 -0.676 -0.310 -0.420 0.367 -0.205 -0.092 -0.121 0.113
1min 2018 -0.512 -0.269 -0.358 0.243 -0.171 -0.086 -0.109 0.085
1min 2019 -0.453 -0.218 -0.324 0.235 -0.128 0.025 -0.049 0.153
1min 2020 -0.493 -0.311 -0.384 0.182 -0.230 -0.010 -0.101 0.220
1min 2021 -0.435 -0.292 -0.356 0.143 -0.164 -0.006 -0.068 0.158
1min 2022 -0.428 -0.244 -0.328 0.184 -0.175 -0.079 -0.106 0.096
1min 2023 -0.596 -0.245 -0.382 0.351 -0.173 -0.025 -0.079 0.148
1min 2024 -0.512 -0.277 -0.385 0.235 -0.149 0.060 -0.019 0.209
5min 2017 -0.452 -0.278 -0.363 0.174 -0.271 -0.172 -0.197 0.099
5min 2018 -0.359 -0.187 -0.277 0.172 -0.236 -0.010 -0.088 0.227
5min 2019 -0.531 -0.187 -0.347 0.344 -0.226 0.070 -0.043 0.296
5min 2020 -0.296 -0.255 -0.319 0.041 -0.304 0.063 -0.059 0.367
5min 2021 -0.309 -0.275 -0.308 0.034 0.039 0.096 0.061 0.057
5min 2022 -0.191 -0.039 -0.107 0.151 -0.197 -0.075 -0.119 0.122
5min 2023 -0.421 -0.208 -0.312 0.213 -0.148 -0.025 -0.069 0.123
5min 2024 -0.513 -0.219 -0.352 0.294 -0.177 -0.066 -0.121 0.111
10min 2017 -0.323 -0.167 -0.270 0.156 -0.159 0.008 -0.055 0.167
10min 2018 -0.482 -0.165 -0.306 0.318 -0.257 -0.002 -0.133 0.255
10min 2019 -0.393 -0.156 -0.289 0.237 -0.249 0.066 -0.066 0.315
10min 2020 -0.199 -0.215 -0.256 0.016 -0.241 0.118 -0.034 0.359
10min 2021 -0.354 -0.254 -0.309 0.100 -0.166 -0.045 -0.086 0.121
10min 2024 -0.522 -0.243 -0.378 0.279 -0.245 -0.012 -0.082 0.233
15min 2018 -0.371 -0.154 -0.263 0.217 -0.158 0.014 -0.058 0.172
15min 2019 -0.463 -0.105 -0.267 0.358 -0.299 0.051 -0.075 0.350
15min 2020 -0.331 -0.211 -0.275 0.120 -0.319 0.194 -0.014 0.513
15min 2021 -0.313 -0.240 -0.282 0.073 -0.215 0.113 -0.017 0.328
15min 2024 -0.583 -0.212 -0.380 0.371 -0.264 -0.032 -0.117 0.232
Appendix B: Supplementary Figures
Figure B.1: Quantile–Quantile plots of Bitcoin log-returns in 2019 at multiple sampling frequencies (1-minute, 5-minute, 10-minute, and 15-minute
(a) 1-minute
(b) 5-minute
(c) 10-minute
(d) 15-minute
Figure B.2: Quantile–Quantile plots of Bitcoin log-returns in 2021 at multiple sampling frequencies (1-minute, 5-minute, 10-minute, and 15-minute
(a) 1-minute
(b) 5-minute
(c) 10-minute
(d) 15-minute
Figure B.3: Quantile–Quantile plots of Bitcoin log-returns in 2024 at multiple sampling frequencies (1-minute, 5-minute, 10-minute, and 15-minute
(a) 1-minute
(b) 5-minute
(c) 10-minute
(d) 15-minute
Figure B.4: Z-score histograms of Bitcoin log-returns in 2019 at multiple sampling frequencies (1-minute, 5-minute, 10-minute, and 15-minute), with densities plotted on a logarithmic scale to visualise tail behaviour
(a) 1-minute
(b) 5-minute
(c) 10-minute
(d) 15-minute
Figure B.5: Z-score histograms of Bitcoin log-returns in 20121 at multiple sampling frequencies (1-minute, 5-minute, 10-minute, and 15-minute), with densities plotted on a logarithmic scale to visualise tail behaviour
(a) 1-minute
(b) 5-minute
(c) 10-minute
(d) 15-minute
Figure B.6: Z-score histograms of Bitcoin log-returns in 2024 at multiple sampling frequencies (1-minute, 5-minute, 10-minute, and 15-minute), with densities plotted on a logarithmic scale to visualise tail behaviour
(a) 1-minute
(b) 5-minute
(c) 10-minute
(d) 15-minute
Appendix C: Code and Data Availability
All code used for data processing, analysis, and figure generation is available from the authors upon request. The Bitcoin high-frequency price data are publicly accessible from Kaggle at https://www.kaggle.com/datasets/mczielinski/bitcoin-historical-data.

### v2

@@
 Master’s Student, MAGEFI – Magistère in International Economics and Finance University of Bordeaux, France
-July 1, 2025
+September 20, 2025
 We assess the applicability of rough volatility models to Bitcoin realised volatility using the normalised $p$-variation framework of Cont and Das, (2024).
@@
 Using longer or multiple windows risked including structural changes or requiring imputation for extensive gaps.
-Let $T=129{,}600$ minutes denote the target window length.
+Let T=129,600T=129{,}600 minutes denote the target window length.
 The search space was defined as the set of all candidate windows of length $T$, starting at time $t_{s}$ and ending at $t_{e}=t_{s}+T-1$, with $t_{s}$ incremented in daily steps (1,440 minutes).
 For each candidate window, we constructed the corresponding sequence of timestamps $\{t_{i}\}_{i=1}^{T}$ and computed the number of missing Close price observations $m(t_{s})$.
-The optimal window $[t_{s}^{*},t_{e}^{*}]$ was selected as:
+The optimal window $t_{s}^{*},t_{e}^{*}]$ was selected as:
 $(t_{s}^{*},t_{e}^{*})=\arg\min_{t_{s}}m(t_{s})$
@@
 This follows the approach of Gatheral et al., (2018) and Cont and Das, (2024), who focus on volatility time series for roughness estimation.
+A potential concern, highlighted in Fukasawa et al., (2022), is that realized volatility constructed from single low-frequency returns
+$RV^{\Delta}_{t}=|r_{t,\Delta}|$
+may be contaminated by white-noise components, leading to artificially rough behaviour.
+To address this, we also consider a noise-robust construction: for each frequency $\Delta$, one realized volatility observation is computed as
+$\widetilde{RV}^{\Delta}_{t}=\left(\sum_{i=1}^{\Delta}r_{t,i}^{2}\right)^{1/2},$
+where $\{r_{t,i}\}$ are one-minute log-returns within the $\Delta$-minute interval.
+This definition ensures that each volatility estimate incorporates multiple high-frequency increments, thereby mitigating the influence of white-noise domination.
+As reported in Appendix A.12, the use of $\widetilde{RV}^{\Delta}_{t}$ does not alter our main findings: the statistic $W(p)$ remains strictly negative across all years and frequencies, indicating that the absence of a zero-crossing cannot be explained by noise contamination alone.
 3.4 Diagnostic Exploration: Roughness vs $K$
@@
 3.5 Final Estimation Setup
-[…]
+For each $(\text{year},\text{frequency})$ pair, we use the optimal block parameters $(K,n)$ determined in Section 3.4.
+The block sums $S_{j}$ are computed, and the normalised $p$-variation statistic is evaluated as:
 $W(L,K,p)=\frac{1}{K}\sum_{j=1}^{K}|S_{j}|^{p}.$ (8)
@@
 As shown in Table 3, $\log W(L,K,p)$ remains strictly negative across all years, frequencies, and both the standard and extended $p$-grids.
-No zero-crossing is observed in any configuration, with maximum values not exceeding $-0.009$, even under the wide grid.
+No zero-crossing is observed in any configuration, with maximum values not exceeding 0.009-0.009, even under the wide grid.
 This behaviour deviates from the expected pattern under homogeneous roughness, where a unique zero-crossing at $p=\hat{p}$ would yield a well-defined roughness exponent $H=1/\hat{p}$.
@@
 15-min 2017, 2022, 2023
-Table A.2: Summary of missing data and large gaps ($>$1 hour) by frequency and year in retained 90-day BTC windows.
-Frequency Year Total Missing Minutes Gaps $>$1 Hour (Count)
+Table A.2: Summary of missing data and large gaps (>>1 hour) by frequency and year in retained 90-day BTC windows.
+Frequency Year Total Missing Minutes Gaps >>1 Hour (Count)
 1-min 2017 1586 0
@@
 Table A.5: Structural Breakpoints Detected by Binary Segmentation (Maximum 5 per Series)
-Frequency Year $N_{\text{obs}}$Breaks Break Indices
+Frequency Year $N_{\text{obs}}$ #Breaks Break Indices
 1min 2017 $128,013$ 5 [81,225, 93,345, 98,420, 110,855, 118,520]
@@
 1min 2020 -2.964 0.007 -0.555 2.971 -0.001 0.554 0.108 0.555
-1min 2021 -0.064 1.407 0.229 1.471 -0.036 $0$0.108 0.036
+1min 2021 -0.064 1.407 0.229 1.471 -0.036 0.000 -0.108 0.036
 1min 2022 -2.379 0.005 -0.654 2.384 0.000 0.593 0.117 0.593
@@
 5min 2018 -0.014 0.003 -0.123 0.017 -0.005 0.242 0.047 0.247
-5min 2019 -0.105 0.533 0.130 0.638 -0.002 $0$0.073 0.002
+5min 2019 -0.105 0.533 0.130 0.638 -0.002 0.000 -0.073 0.002
 5min 2020 -2.053 0.009 -0.487 2.062 -0.000 1.004 0.293 1.004
@@
 5min 2024 -0.041 0.007 -0.406 0.048 -0.010 -0.312 -0.063 0.302
-10min 2017 -0.084 0.068 0.008 0.152 -0.949 $0$0.333 0.949
-10min 2018 -0.016 0.135 0.029 0.151 -0.115 $0$0.045 0.115
+10min 2017 -0.084 0.068 0.008 0.152 -0.949 0.000 -0.333 0.949
+10min 2018 -0.016 0.135 0.029 0.151 -0.115 0.000 -0.045 0.115
 10min 2019 -0.360 0.002 -0.079 0.362 -0.168 0.001 -0.048 0.168
@@
 1min 2018 -4.314 0.049 -0.849 4.363 -0.002 1.272 0.328 1.274
-1min 2019 -3.714 0.082 -0.754 3.796 -0.002 $0$0.096 0.002
+1min 2019 -3.714 0.082 -0.754 3.796 -0.002 0.000 -0.096 0.002
 1min 2020 -2.628 0.122 -0.451 2.750 -0.000 0.290 0.073 0.290
@@
 5min 2017 -3.164 0.101 -0.629 3.265 -0.000 1.204 0.339 1.204
-5min 2018 -0.279 0.057 -0.063 0.336 -0.010 -$0$0.071 0.010
+5min 2018 -0.279 0.057 -0.063 0.336 -0.010 -0.000 -0.071 0.010
 5min 2019 -0.014 1.010 0.281 1.025 -0.015 0.125 0.026 0.140
@@
 5min 2023 -1.919 0.066 -0.311 1.985 -0.001 0.367 0.068 0.368
-5min 2024 -1.503 0.079 -0.202 1.582 -0.007 -$0$0.130 0.007
-10min 2017 -0.006 0.131 0.045 0.137 -0.793 $0$0.195 0.793
-10min 2018 -0.121 0.049 -0.022 0.170 -0.561 $0$0.130 0.561
+5min 2024 -1.503 0.079 -0.202 1.582 -0.007 -0.000 -0.130 0.007
+10min 2017 -0.006 0.131 0.045 0.137 -0.793 0.000 -0.195 0.793
+10min 2018 -0.121 0.049 -0.022 0.170 -0.561 0.000 -0.130 0.561
 10min 2019 -0.024 0.063 0.018 0.087 0.000 0.500 0.133 0.500
@@
 15min 2024 -0.583 -0.212 -0.380 0.371 -0.264 -0.032 -0.117 0.232
+Table A.12: Minimum and maximum values of $\log W(L,K,p)$ for each (frequency, year) pair, using the noise-robust realised volatility $\widetilde{RV}^{\Delta}_{t}$.
+Frequency Year Standard Grid Wide Grid
+Min Max Min Max
+1min 2017 -1.797 -0.090 -1.797 -0.009
+1min 2018 -3.454 -0.110 -3.454 -0.011
+1min 2019 -3.149 -0.114 -3.149 -0.011
+1min 2020 -1.857 -0.124 -1.857 -0.012
+1min 2021 -2.377 -0.098 -2.377 -0.010
+1min 2022 -5.021 -0.154 -5.021 -0.015
+1min 2023 -6.177 -0.199 -6.177 -0.020
+1min 2024 -5.249 -0.175 -5.249 -0.018
+5min 2017 -1.188 -0.067 -1.188 -0.007
+5min 2018 -2.738 -0.089 -2.738 -0.009
+5min 2019 -2.625 -0.093 -2.625 -0.009
+5min 2020 -1.434 -0.104 -1.434 -0.010
+5min 2021 -1.855 -0.079 -1.855 -0.008
+5min 2022 -4.442 -0.133 -4.442 -0.013
+5min 2023 -5.658 -0.176 -5.658 -0.018
+5min 2024 -4.778 -0.155 -4.778 -0.016
+10min 2017 -1.093 -0.062 -1.093 -0.006
+10min 2018 -2.628 -0.084 -2.628 -0.008
+10min 2019 -2.542 -0.089 -2.542 -0.009
+10min 2020 -1.388 -0.100 -1.388 -0.010
+10min 2021 -1.782 -0.075 -1.782 -0.008
+10min 2024 -4.831 -0.150 -4.831 -0.015
+15min 2018 -2.588 -0.083 -2.588 -0.008
+15min 2019 -2.446 -0.087 -2.446 -0.009
+15min 2020 -1.402 -0.099 -1.402 -0.010
+15min 2021 -1.908 -0.073 -1.908 -0.007
+15min 2024 -4.807 -0.150 -4.807 -0.015
 Appendix B: Supplementary Figures

### v3

@@
-Scale-Dependent Multifractality in Bitcoin Realised Volatility: Implications for Rough Volatility Modelling
-Scale-Dependent Multifractality in Bitcoin Realised Volatility: Implications for Rough Volatility Modelling
+Multifractality in Bitcoin Realized Volatility: Implications for Rough Volatility Modelling
+Multifractality in Bitcoin Realized Volatility: Implications for Rough Volatility Modelling
 Milan Pontiggia
 Master’s Student, MAGEFI – Magistère in International Economics and Finance University of Bordeaux, France
-September 20, 2025
-We assess the applicability of rough volatility models to Bitcoin realised volatility using the normalised $p$-variation framework of Cont and Das, (2024).
-Applying this model-free estimator to high-frequency Bitcoin data from 2017 to 2024 across multiple sampling resolutions, we find that the normalised statistic remains strictly negative throughout, precluding the estimation of a valid roughness index.
+September 28, 2025
+We assess the applicability of rough volatility models to Bitcoin realized volatility using the normalised $p$-variation framework of Cont and Das, (2024).
+Applying this model-free estimator to high-frequency Bitcoin data from 2017 to 2024 across multiple sampling resolutions, we find that the normalised statistic remains strictly negative, precluding the estimation of a valid roughness index.
 Stationarity tests and robustness checks reveal no significant evidence of non-stationarity or structural breaks as explanatory factors.
 Instead, convergent evidence from three complementary diagnostics, namely Multifractal Detrended Fluctuation Analysis, log-log moment scaling, and wavelet leaders, reveals a multifractal structure in Bitcoin volatility.
-This scale-dependent behaviour violates the homogeneity assumptions underlying rough volatility estimation and accounts for the estimator’s systematic failure.
+This behaviour violates the homogeneity assumptions underlying rough volatility estimation and accounts for the estimator’s systematic failure.
 These findings suggest that while rough volatility models perform well in traditional markets, they are structurally misaligned with the empirical features of Bitcoin volatility.
@@
 The rough volatility framework has emerged as a prominent approach for modelling the fine-scale dynamics of financial market volatility.
-This development is grounded in empirical findings that indicate that realised volatility paths possess low Hölder regularity, which implies irregularity in their sample paths at high frequencies.
+This development is grounded in empirical findings that indicate that realized volatility paths possess low Hölder regularity, which implies irregularity in their sample paths at high frequencies.
 Rough stochastic volatility models have been shown to replicate empirical features commonly observed in financial markets, including volatility clustering, short-term persistence, and the skew and term structure of implied volatility surfaces in equity and index markets Gatheral et al., (2018); Fukasawa et al., (2022).
@@
 Unlike traditional equity or FX markets, Bitcoin exhibits fragmented liquidity across exchanges Makarov and Schoar, (2020), frequent discontinuities due to jumps Lahmiri and Bekiros, (2020), regime-switching volatility Caporale and Zekokh, (2021), and an evolving market microstructure shaped by protocol changes and shifting incentives Easley et al., (2019).
-This motivates a closer empirical examination of whether rough volatility assumptions hold in such a market.
+This motivates an empirical examination of whether rough volatility assumptions hold in such a market.
 In this context, the term roughness refers to the local irregularity of the volatility path, quantified in terms of its Hölder regularity.
@@
 Prior applications of rough volatility methodologies to Bitcoin have predominantly employed parametric inference techniques or heuristic scaling diagnostics, often without systematically assessing assumptions of homogeneous Hölder continuity Bianchi et al., (2022).
-Empirical studies on cryptocurrencies more broadly, including Liu and Tsyvinski, (2021), have concentrated on asset pricing and factor models, without directly addressing the scaling properties of volatility.
 The normalised $p$-variation estimator recently introduced by Cont and Das, (2024) provides a model-free approach to estimating the pathwise roughness of volatility signals.
@@
 These properties make it especially suitable for high-frequency volatility data in contexts like cryptocurrency markets, where noise and irregular sampling are common.
-This study builds on the normalised $p$-variation estimator introduced by Cont and Das, (2024) to perform a systematic evaluation of the roughness of Bitcoin realised volatility across a range of sampling frequencies and time horizons.
+This study builds on the normalised $p$-variation estimator introduced by Cont and Das, (2024) to perform a systematic evaluation of the roughness of Bitcoin realized volatility across a range of sampling frequencies and time horizons.
 This analysis examines the applicability of rough volatility models to Bitcoin and contributes to a deeper understanding of the structural features that differentiate cryptocurrency volatility dynamics.
@@
 However, the applicability of such models to structurally distinct markets such as cryptocurrencies remains largely unexplored.
-Bitcoin’s volatility displays structural features such as multiscale clustering, episodic jumps, and irregular liquidity that differ markedly from the conditions under which rough volatility models have been validated.For instance, Bianchi et al.
-(2022) report low Hurst exponents for Bitcoin by fitting rough stochastic volatility models and performing scaling regressions on realised volatility.
+Bitcoin’s volatility displays structural features such as multiscale clustering, episodic jumps, and irregular liquidity that differ markedly from the conditions under which rough volatility models have been validated.For instance, Bianchi et al., (2022) report low Hurst exponents for Bitcoin by fitting rough stochastic volatility models and performing scaling regressions on realized volatility.
 However, their approach implicitly assumes a global scaling regime and stationarity, without subjecting these conditions to empirical verification.
 These assumptions are critical, and their violation under multifractal dynamics or structural irregularities undermines the reliability of such roughness estimates.
-Crucially, no prior work has systematically examined whether Bitcoin realised volatility meets the homogeneity and stationarity conditions required for consistent estimation via the normalised $p$-variation method.
-Moreover, recent research indicates that the observed roughness of realised volatility may arise from estimation error rather than genuine features of the spot volatility process Cont and Das, (2024); Rogers, (2019).
+Crucially, no prior work has systematically examined whether Bitcoin realized volatility meets the homogeneity and stationarity conditions required for consistent estimation via the normalised $p$-variation method.
+Moreover, recent research indicates that the observed roughness of realized volatility may arise from estimation error rather than genuine features of the spot volatility process Cont and Das, (2024); Rogers, (2019).
 In the case of Bitcoin, where multifractality and non-stationarity are theoretically plausible features, this raises serious doubts about the robustness and interpretive value of roughness estimates.
-To our knowledge, no prior empirical study has systematically applied the $p$-variation diagnostic to Bitcoin realised volatility, nor has tested its structural compatibility with the assumptions underlying rough volatility models.
+To our knowledge, no prior empirical study has systematically applied the $p$-variation diagnostic to Bitcoin realized volatility, nor has tested its structural compatibility with the assumptions underlying rough volatility models.
 It remains empirically unestablished whether Bitcoin volatility admits a well-defined roughness exponent.
@@
 1.3 Research Question and Empirical Puzzle
-This study tests whether Bitcoin realised volatility satisfies the homogeneous scaling assumptions necessary for consistent roughness estimation under the rough volatility paradigm.
-We investigate this question by applying the normalised $p$-variation methodology of Cont and Das, (2024) to Bitcoin realised volatility across multiple years and sampling frequencies.
+This study tests whether Bitcoin realized volatility satisfies the homogeneous scaling assumptions necessary for roughness estimation under the rough volatility paradigm.
+We investigate this question by applying the normalised $p$-variation methodology of Cont and Das, (2024) to Bitcoin realized volatility across multiple years and sampling frequencies.
 Preliminary results indicate that, in contrast to equity and foreign exchange markets, Bitcoin volatility does not appear to admit a well-defined roughness index.
@@
 This study makes three contributions to the literature on volatility modelling and the structural properties of cryptocurrency markets.
-First, it offers an extensive application of the normalised $p$-variation framework to Bitcoin realised volatility across multiple years and sampling frequencies, making the analysis more comprehensive than any prior roughness estimation in cryptocurrency markets.
+First, it offers an extensive application of the normalised $p$-variation framework to Bitcoin realized volatility across multiple years and sampling frequencies, making the analysis more comprehensive than any prior roughness estimation in cryptocurrency markets.
 The analysis reveals a systematic and robust failure of roughness estimation across all empirical configurations.
@@
 This structure is incompatible with the homogeneous scaling assumptions required by rough volatility models.
-Third, the results contribute to the evaluation of the rough volatility paradigm by showing that, while effective in equity and foreign exchange markets, rough stochastic volatility models face structural challenges when applied to Bitcoin realised volatility.
+Third, the results contribute to the evaluation of the rough volatility paradigm by showing that, while effective in equity and foreign exchange markets, rough stochastic volatility models face structural challenges when applied to Bitcoin realized volatility.
 This finding underscores the need for models that explicitly incorporate multifractal and scale-dependent dynamics to capture the unique characteristics of cryptocurrency markets.
@@
 Empirical support for the rough volatility hypothesis has emerged from multiple methodological approaches.
-These include the log-regression of realised volatility increments Gatheral et al., (2018), parametric estimation of fractional stochastic volatility models Fukasawa et al., (2022), and non-parametric diagnostics based on $p$-variation Cont and Das, (2024).
+These include the log-regression of realized volatility increments Gatheral et al., (2018), parametric estimation of fractional stochastic volatility models Fukasawa et al., (2022), and non-parametric diagnostics based on $p$-variation Cont and Das, (2024).
 Across these methods, estimated Hurst exponents in equity and index markets typically fall in the range 0.1 to 0.3.
-These findings have motivated the adoption of rough stochastic volatility models, originally proposed by Gatheral and Rosenbaum and empirically validated by Bennedsen, Lunde, and Pakkanen Bennedsen et al., (2016), in both academic research and quantitative finance practice.
 Rough models have demonstrated empirical success in capturing the fine-scale behaviour of implied volatility surfaces and short-term volatility dynamics.
 The applicability of the rough volatility paradigm across asset classes remains contested.
-Rogers, (2019) and Cont and Das, (2024) have emphasised that low roughness estimates derived from realised volatility may not reflect intrinsic properties of the spot volatility process, but instead result from structural features of the data or estimation artefacts.
-For example, Rogers, (2019) showed that microstructure noise or irregular sampling can generate artificially low Hurst exponents, cautioning against interpreting realised path roughness as direct evidence of true volatility roughness.
+Rogers, (2019) and Cont and Das, (2024) have emphasised that low roughness estimates derived from realized volatility may not reflect intrinsic properties of the spot volatility process, but instead result from structural features of the data or estimation artefacts.
+For example, Rogers, (2019) showed that microstructure noise or irregular sampling can generate artificially low Hurst exponents, cautioning against interpreting realized path roughness as direct evidence of true volatility roughness.
 These confounding effects, including multifractal scaling, may induce apparent roughness even when the underlying volatility process is not genuinely rough.
@@
 This framework provides a model-free, pathwise-consistent estimator of roughness based on the scaling behaviour of the normalised $p$-variation statistic $W(L,K,p)$, computed over a sequence of block partitions of the observed time series.
-Intuitively, $W(L,K,p)$ measures how the $p$-th variation of the process accumulates across blocks, normalised in such a way that for a self-similar process with a global scaling law, the statistic equals one exactly at the correct scaling exponent $p^{*}$.
+$W(L,K,p)$ measures how the $p$-th variation of the process accumulates across blocks, normalised in such a way that for a self-similar process with a global scaling law, the statistic equals one exactly at the correct scaling exponent $p^{*}$.
 Thus, plotting $\log W(L,K,p)$ as a function of $1/p$, one expects a zero-crossing at a unique value $p=p^{*}$, which defines the variation index and yields the corresponding roughness index $H=1/p^{*}$.
@@
 For reference, the formal definition of $W(L,K,p)$ is provided in Section 3.1.
-The normalised $p$-variation estimator is non-parametric, robust to non-Gaussian features, and directly applicable to realised volatility series.
+The normalised $p$-variation estimator is non-parametric, robust to non-Gaussian features, and directly applicable to realized volatility series.
 It serves not only as a roughness estimator but also as a structural diagnostic of the underlying process.
@@
 Recent applications of the method to equity markets have validated its efficacy Cont and Das, (2024), producing roughness estimates consistent with the existing literature.
-In particular, Cont and Das apply the normalised $p$-variation estimator to 5-minute realised volatility of the S&P 500 using the Oxford-Man Institute dataset.
+In particular, Cont and Das apply the normalised $p$-variation estimator to 5-minute realized volatility of the S&P 500 using the Oxford-Man Institute dataset.
 They observe a well-defined zero-crossing in $\log W(L,K,p)$, with the resulting roughness index estimated in the range $H\in[0.05,0.25]$, consistent with prior findings from log-regression methods.
@@
 High-frequency Bitcoin returns and trading volume exhibit multifractal scaling when analysed using Multifractal Detrended Fluctuation Analysis Mensi et al., (2019); Matteo, (2007), and complex multiscale dependence has been observed in cross-correlation structures Watorek et al., (2022).
-Additional research shows that Bitcoin realised volatility also displays a multifractal structure, though weaker than that of returns Zhou et al., (2022).
+Additional research shows that Bitcoin realized volatility also displays a multifractal structure, though weaker than that of returns Zhou et al., (2022).
 These findings support the view that multifractality is a robust and persistent feature of financial markets and offer a plausible explanation for the structural deviations from monofractal scaling observed in Bitcoin volatility.
@@
 15-min 7,927 8,640 91.75
-A 90% completeness threshold was imposed at each frequency to balance data quality with temporal coverage.
+A 90% completeness threshold was imposed at each frequency.
 For 1-minute and 5-minute data, all years from 2017 to 2024 met this criterion.
@@
 Large temporal gaps, defined as sequences of missing data exceeding one hour, were systematically identified at each frequency.
-Such gaps were rare for 1-minute and 5-minute data, but more common at coarser resolutions, especially in early years and in 2024.
 Full gap statistics are provided in Appendix Table A.2.
@@
 No filtering was applied beyond the initial screening, all flagged returns were retained, and no values were removed or modified.
-3.3 Realised Volatility Construction
+3.3 Realized Volatility Construction
 For each year and frequency configuration, we define a proxy for high-frequency volatility by taking the modulus of log-returns at the original sampling resolution.
@@
 This absolute return series preserves the temporal resolution of the data and captures the local magnitude of return fluctuations.
-For notational convenience, we refer to $\{RV_{t}\}$ as the realised volatility series, although it is more accurately a proxy for instantaneous volatility rather than an integrated measure.
+For notational convenience, we refer to $\{RV_{t}\}$ as the realized volatility series, although it is more accurately a proxy for instantaneous volatility rather than an integrated measure.
 Formally, for sampling frequency $f$ and year $Y$, let $X(f,Y)=\{r_{t}(f,Y)\}_{t=1}^{L_{f,Y}}$ denote the log-return series. We construct
@@
 Prior to finalising the roughness estimation procedure, we conduct a diagnostic analysis to examine how the estimated roughness $H_{L,K}$ depends on the number of blocks $K$ (which controls the partition size).
-This step ensures we choose an appropriate $K$ for each case, as the estimate can be sensitive to this choice.
 The diagnostic analysis evaluates the behaviour of the roughness estimator $\widehat{H}^{L,K}$ as a function of the number of blocks $K$.
@@
 To mitigate these issues, Cont and Das, (2024) recommend selecting $K$ such that $K\approx\sqrt{L}$, where $L$ is the number of available observations.
-This choice ensures a trade-off between bias and variance, grounded in the theoretical convergence properties of the normalised $p$-variation statistic:
+This choice ensures a trade-off between bias and variance:
 $\lim_{K\to\infty}\lim_{L\to\infty}W(L,K,p)\to w(x,p,\pi)$ (7)
 and maintains a sufficient block size $n=L/K$ to stabilise the estimation.
-Following the recommendation of Cont and Das, (2024), we set $K_{\text{opt}}=\lfloor\sqrt{N}\rfloor$, where $N$ is the number of realised volatility observations for the given year and frequency.
+Following the recommendation of Cont and Das, (2024), we set $K_{\text{opt}}=\lfloor\sqrt{N}\rfloor$, where $N$ is the number of realized volatility observations for the given year and frequency.
 The corresponding block size is set to $n_{\text{opt}}=K_{\text{opt}}$, ensuring an equal partition.
@@
 The presence of a clear zero-crossing is a necessary condition for Bitcoin’s volatility to follow a rough volatility scaling law (Hypothesis H1).
-Following Cont and Das, (2024), we apply the estimator directly to the series of realised volatility magnitudes, defined as $RV_{t}=|r_{t}|$, where $r_{t}$ is the log-return at time $t$.
+Following Cont and Das, (2024), we apply the estimator directly to the series of realized volatility magnitudes, defined as $RV_{t}=|r_{t}|$, where $r_{t}$ is the log-return at time $t$.
 This approach avoids any transformation to log-volatility or its increments, and the resulting roughness index $\hat{H}$ captures the fine-scale regularity of the volatility path itself.
@@
 3.6.1 Non-Stationarity Diagnostics
-To assess whether the failure of the $p$-variation estimator may stem from non-stationarity or structural breaks, we apply three complementary diagnostics to each realised volatility series: the Augmented Dickey-Fuller (ADF) test for mean stationarity, rolling window analysis to evaluate local stability of mean and variance, and structural break detection to identify discrete regime shifts.
+To assess whether the failure of the $p$-variation estimator may stem from non-stationarity or structural breaks, we apply three complementary diagnostics to each realized volatility series: the Augmented Dickey-Fuller (ADF) test for mean stationarity, rolling window analysis to evaluate local stability of mean and variance, and structural break detection to identify discrete regime shifts.
 Together, these methods provide a comprehensive assessment of both gradual and abrupt departures from stationarity.
@@
 3.6.2 Multifractality Diagnostics
-To assess whether Bitcoin’s realised volatility exhibits a multifractal temporal structure, we apply three complementary diagnostics: Multifractal Detrended Fluctuation Analysis (MF-DFA), log-log moment scaling, and wavelet leaders analysis.
-Each method is implemented systematically across all (year, frequency) combinations and applied uniformly to both log-return and realised volatility series.
+To assess whether Bitcoin’s realized volatility exhibits a multifractal temporal structure, we apply three complementary diagnostics: Multifractal Detrended Fluctuation Analysis (MF-DFA), log-log moment scaling, and wavelet leaders analysis.
+Each method is implemented systematically across all (year, frequency) combinations and applied uniformly to both log-return and realized volatility series.
 This convention is maintained throughout and not repeated in the descriptions that follow.
@@
 is computed, where $F_{q}(s)$ denotes the $q$th-order fluctuation function at scale $s$.
-Summary statistics for the resulting $H(q)$ spectra (including minimum, maximum, mean, and spectral width) are reported in Appendix Table A.6 (log-returns) and Table A.9 (realised volatility).
+Summary statistics for the resulting $H(q)$ spectra (including minimum, maximum, mean, and spectral width) are reported in Appendix Table A.6 (log-returns) and Table A.9 (realized volatility).
 Log-log moment scaling analysis is used to estimate empirical scaling exponents $\zeta_{q}$ by computing, for each moment order $q\in[-4,4]$, the slope of the scaling relation
@@
 This procedure is implemented over a dense grid of lags $\tau$.
-Complete results are presented in Appendix Table A.7 (log-returns) and Table A.10 (realised volatility).
+Complete results are presented in Appendix Table A.7 (log-returns) and Table A.10 (realized volatility).
 Wavelet leaders multifractal analysis estimates the scaling relation for each moment order $q\in[-4,4]$, the scaling relation
@@
 Wavelet leaders, constructed from the local suprema of wavelet coefficients across scales, offer robustness to non-stationarities and facilitate the isolation of multiscale fluctuations from deterministic trends.
-Full results are presented in Appendix Table A.8 (log-returns) and Table A.11 (realised volatility).
+Full results are presented in Appendix Table A.8 (log-returns) and Table A.11 (realized volatility).
 A non-linear relationship between $H(q)$ or $\zeta_{q}$ and the moment order $q$ is a hallmark of multifractality, indicating scale-dependent dynamics.
@@
 As detailed in Section 3.4, examining the behaviour of $\widehat{H}^{L,K}$ across varying $K$ offers diagnostic insight into scaling stability and guides the selection of a robust block size.
-Figure 1: Estimated roughness index $\widehat{H}^{L,K}$ versus number of blocks $K$ for BTC realised volatility in 2021 at 1-minute (top-left), 5-minute (top-right), 10-minute (bottom-left), and 15-minute (bottom-right) frequencies.
+Figure 1: Estimated roughness index $\widehat{H}^{L,K}$ versus number of blocks $K$ for BTC realized volatility in 2021 at 1-minute (top-left), 5-minute (top-right), 10-minute (bottom-left), and 15-minute (bottom-right) frequencies.
 As shown in Figure 1, for the 1-minute data, $\widehat{H}^{L,K}$ is clearly positive for small $K$ (up to approximately $K\approx 200$) and then declines sharply, becoming negative for larger $K$.
@@
 Diagnostic results do not support non-stationarity or structural breaks as plausible explanations for the failure of the normalised $p$-variation estimator.
-Bitcoin realised volatility appears stationary with stable local dynamics.
+Bitcoin realized volatility appears stationary with stable local dynamics.
 If non-stationarity were the cause, one would expect to observe systematic trends, level shifts, or persistent regime changes, none of which are observed in the diagnostic results.
@@
 Across all original series, the estimated $H(q)$ spectra exhibit pronounced non-linearity with respect to $q$, indicating scale-dependent behaviour characteristic of multifractality.
-For realised volatility, $H(q)$ remain consistently below 1 across all years and sampling frequencies, typically centred around 0.9 for moderate $q$, with spectral widths between 0.06 and 0.42.
+For realized volatility, $H(q)$ remain consistently below 1 across all years and sampling frequencies, typically centred around 0.9 for moderate $q$, with spectral widths between 0.06 and 0.42.
 Log-returns display lower mean exponents but retain comparable curvature in the scaling function.
@@
 We next apply the log-log moment scaling analysis, an alternative multifractality diagnostic based on the scaling behaviour of $\zeta_{q}$, as described in Section 3.6.2.
-Estimated $\zeta_{q}$ spectra for realised volatility and log-returns are presented in Appendix Tables A.7 and A.10.
-For realised volatility, the $\zeta_{q}$ spectra exhibit pronounced curvature in $q$, with widths ranging from 0.04 to over 2.0 and mean $R^{2}$ values generally exceeding 0.85, indicating stable scaling across temporal resolutions.
+Estimated $\zeta_{q}$ spectra for realized volatility and log-returns are presented in Appendix Tables A.7 and A.10.
+For realized volatility, the $\zeta_{q}$ spectra exhibit pronounced curvature in $q$, with widths ranging from 0.04 to over 2.0 and mean $R^{2}$ values generally exceeding 0.85, indicating stable scaling across temporal resolutions.
 Log-return series show more heterogeneity (especially at lower sampling frequencies) but maintain clear non-linearity in the $\zeta_{q}$ profile for moderate to large $q$.
@@
 The wavelet leaders multifractal analysis offers a robust characterisation of scaling behaviour, particularly in the presence of trends or non-stationarities.
-Estimated $\zeta_{q}$ exponents for both realised volatility and log-return series are presented in Appendix Tables A.8 and A.11.
-For realised volatility, the $\zeta_{q}$ spectra exhibit consistent curvature with spectral widths between 0.04 and 0.51 and mean $R^{2}$ values typically exceeding 0.85.
+Estimated $\zeta_{q}$ exponents for both realized volatility and log-return series are presented in Appendix Tables A.8 and A.11.
+For realized volatility, the $\zeta_{q}$ spectra exhibit consistent curvature with spectral widths between 0.04 and 0.51 and mean $R^{2}$ values typically exceeding 0.85.
 Log-return series show similar but generally weaker curvature, with greater sensitivity to sampling resolution.
 In the shuffled series, spectral widths decline substantially (often by more than 50%) and curvature is notably reduced.
-The MF-DFA, log-log moment scaling, and wavelet leaders analyses consistently indicate that Bitcoin realised volatility exhibits a multifractal temporal structure.
+The MF-DFA, log-log moment scaling, and wavelet leaders analyses consistently indicate that Bitcoin realized volatility exhibits a multifractal temporal structure.
 This conclusion is supported by the marked reduction in spectral curvature and width observed in shuffled controls, which preserve the marginal distribution but remove temporal dependence.
@@
 H1. Bitcoin volatility does not admit a global scaling law compatible with the normalised $p$-variation framework:
-Our primary empirical result, the strict negativity of $\log W(L,K,p)$ across all years, frequencies, and $p$-grids, demonstrates the absence of any stable scaling law compatible with the normalised $p$-variation estimator in Bitcoin realised volatility.
+Our primary empirical result, the strict negativity of $\log W(L,K,p)$ across all years, frequencies, and $p$-grids, demonstrates the absence of any stable scaling law compatible with the normalised $p$-variation estimator in Bitcoin realized volatility.
 This supports the interpretation that the apparent roughness reported in prior literature does not reflect an intrinsic property of the volatility process, but rather the failure of the $p$-variation framework to capture consistent scaling behaviour in this context.
@@
 Our suite of stationarity diagnostics, including the Augmented Dickey-Fuller test for unit roots, rolling mean and variance statistics for local stability, and structural break detection via binary segmentation with an $L_{2}$-based cost function, yields no evidence that non-stationarity or persistent structural instability accounts for the failure of the $p$-variation estimator.
-The realised volatility series are stationary over the studied periods, and the detected breakpoints correspond to local volatility clustering rather than structural regime changes.
+The realized volatility series are stationary over the studied periods, and the detected breakpoints correspond to local volatility clustering rather than structural regime changes.
 We therefore reject $H_{3}$ as a primary explanation for the observed behaviour of $\log W(L,K,p)$, given the absence of non-stationarity or structural instability.
@@
 All three diagnostic approaches, MF-DFA, log-log moment scaling, and wavelet leaders analysis, exhibit consistent non-linear scaling in the original series and show substantial collapse of multifractal indicators under shuffle-based controls.
-This confirms that the observed multifractality in Bitcoin volatility, particularly in realised volatility, arises from genuine temporal correlations rather than from heavy-tailed marginal distributions.
+This confirms that the observed multifractality in Bitcoin volatility, particularly in realized volatility, arises from genuine temporal correlations rather than from heavy-tailed marginal distributions.
 The presence of this multifractal temporal structure implies non-linear, scale-dependent behaviour that directly violates the homogeneity assumption of the normalised $p$-variation framework.
@@
 5.1 Structural Interpretation
-Our analysis demonstrates that Bitcoin realised volatility exhibits pronounced scale dependence and temporal heterogeneity, in contrast to the patterns typically observed in traditional financial assets.
+Our analysis demonstrates that Bitcoin realized volatility exhibits pronounced scale dependence and temporal heterogeneity, in contrast to the patterns typically observed in traditional financial assets.
 Evidence from multiple diagnostics, reinforced by shuffled controls, consistently indicates the presence of multifractality and the absence of a global scaling regime.
@@
 While this study highlights the limitations of rough volatility estimation for Bitcoin and confirms the presence of multifractal structure, several limitations should be acknowledged.
-The study focuses exclusively on realised volatility derived from high-frequency price data.
+The study focuses exclusively on realized volatility derived from high-frequency price data.
 Although well-suited for $p$-variation analysis, this measure captures only a subset of the latent volatility process.
 The scaling behaviour of alternative measures, such as implied volatility or state-space estimates, has not been examined.
-Extending the analysis to these measures would clarify whether multifractality is an intrinsic property of Bitcoin volatility or an artefact of realised volatility estimation, and would offer insights into how market expectations reflect multiscale behaviour.
+Extending the analysis to these measures would clarify whether multifractality is an intrinsic property of Bitcoin volatility or an artefact of realized volatility estimation, and would offer insights into how market expectations reflect multiscale behaviour.
 Second, despite careful preprocessing and partitioning following best practices Cont and Das, (2024), our results remain subject to finite-sample limitations.
@@
 Diagnostic tests rule out non-stationarity, structural breaks, and sampling artefacts as plausible explanations.
-Instead, consistent evidence from MF-DFA, log-log moment scaling, and wavelet leaders indicates that Bitcoin realised volatility follows a robust multifractal temporal structure, driven by dynamic correlations rather than marginal distributional effects.
+Instead, consistent evidence from MF-DFA, log-log moment scaling, and wavelet leaders indicates that Bitcoin realized volatility follows a robust multifractal temporal structure, driven by dynamic correlations rather than marginal distributional effects.
 This violates the homogeneity assumptions required by rough volatility models and accounts for the systematic failure of the estimator.
@@
 15min 2024 -0.220 0.081 -0.037 0.301 -0.009 0.041 0.018 0.050
-Table A.9: Summary of MF-DFA Scaling Exponents $H(q)$ for Bitcoin Realised Volatility (Original vs.
+Table A.9: Summary of MF-DFA Scaling Exponents $H(q)$ for Bitcoin Realized Volatility (Original vs.
 Shuffled, 2017–2024).
@@
 15min 2024 0.747 0.942 0.870 0.195 0.478 0.654 0.527 0.176
-Table A.10: Summary of Log-Log Scaling Exponents for Bitcoin Realised Volatility (Original vs.
+Table A.10: Summary of Log-Log Scaling Exponents for Bitcoin Realized Volatility (Original vs.
 Shuffled, 2017–2024).
@@
 15min 2024 -1.249 0.091 -0.193 1.340 -0.002 0.152 0.038 0.154
-Table A.11: Summary of Wavelet Leaders Scaling Exponents for Bitcoin Realised Volatility (Original vs.
+Table A.11: Summary of Wavelet Leaders Scaling Exponents for Bitcoin Realized Volatility (Original vs.
 Shuffled, 2017–2024).
@@
 15min 2024 -0.583 -0.212 -0.380 0.371 -0.264 -0.032 -0.117 0.232
-Table A.12: Minimum and maximum values of $\log W(L,K,p)$ for each (frequency, year) pair, using the noise-robust realised volatility $\widetilde{RV}^{\Delta}_{t}$.
+Table A.12: Minimum and maximum values of $\log W(L,K,p)$ for each (frequency, year) pair, using the noise-robust realized volatility $\widetilde{RV}^{\Delta}_{t}$.
 Frequency Year Standard Grid Wide Grid